### 07 Feb 2024 v1

# by Leveraging CLIP Latent Space

Abstract
* indicates equal contribution, $\dagger$ Corresponding Author: maitreya.patel@asu.edu.
Despite the recent advances in personalized text-to-image (P-T2I) generative models, subject-driven T2I remains challenging. The primary bottlenecks include 1) Intensive training resource requirements, 2) Hyper-parameter sensitivity leading to inconsistent outputs, and 3) Balancing the intricacies of novel visual concept and composition alignment. We start by re-iterating the core philosophy of T2I diffusion models to address the above limitations. Predominantly, contemporary subject-driven T2I approaches hinge on Latent Diffusion Models (LDMs), which facilitate T2I mapping through cross-attention layers. While LDMs offer distinct advantages, P-T2I methods’ reliance on the latent space of these diffusion models significantly escalates resource demands, leading to inconsistent results and necessitating numerous iterations for a single desired image. Recently, ECLIPSE has demonstrated a more resource-efficient pathway for training UnCLIP-based T2I models, circumventing the need for diffusion text-to-image priors. Building on this, we introduce $\lambda$-ECLIPSE [^1] The designation $\lambda$-ECLIPSE is inspired by its conceptual alignment with the $\lambda$-calculus. In this context, the $\lambda$-ECLIPSE model functions similarly to a functional abstraction within $\lambda$-calculus, where it effectively binds variables. These variables, in our case, represent novel visual concepts that are integrated through composition prompts.. Our method illustrates that effective P-T2I does not necessarily depend on the latent space of diffusion models. $\lambda$-ECLIPSE  achieves single, multi-subject, and edge-guided T2I personalization with just 34M parameters and is trained on a mere 74 GPU hours using 1.6M image-text interleaved data. Through extensive experiments, we also establish that $\lambda$-ECLIPSE  surpasses existing baselines in composition alignment while preserving concept alignment performance, even with significantly lower resource utilization.
Figure 1: Our $\lambda$-ECLIPSE  can estimate subject-specific image embeddings while maintaining the balance between concept and composition alignment. $\lambda$-ECLIPSE  supports multi-concept personalization and provides additional controls (i.e., edge map) for image generation. In this work, we show that it is possible to perform subject-driven T2I in the latent space of a pre-trained CLIP model without depending on the diffusion UNet models; effectively reducing the heavy resource requirements.
1 Introduction
The field of text-to-image (T2I) diffusion models has recently witnessed remarkable advancements, achieving greater photorealism and enhanced adherence to textual prompts. This has catalyzed the emergence of diverse applications, notably in controllable T2I models and personalized T2I (P-T2I) models. P-T2I, in particular, encompasses the intricate task of learning and reproducing novel visual concepts or subjects in varied contexts, beyond mere replication of reference images. This personalization task is challenging, necessitating a nuanced balance between concept and compositional alignment. The complexity escalates further when multi-subject personalization is desired.
Early works in this field, such as Textual Inversion [12] and DreamBooth [38], employed per-concept optimization strategies involving fine-tuning certain parameters within T2I diffusion models. However, these methods struggle with generalization challenges and are time-intensive. Contemporary research is pivoting towards fast personalization techniques that involve training hypernetworks and integrating new layers or parameters within pre-trained diffusion UNet models. As summarized in Table 1, despite their relative parameter efficiency, these methods are still resource-intensive (especially for multi-concept personalization), largely due to their reliance on the latent spaces of diffusion models. Additionally, controlling the inherent randomness in diffusion models remains a significant challenge, often resulting in inconsistent outputs and necessitating multiple attempts to generate a single accurate image.
Upon further investigation, we find that most subject-driven T2I approaches build upon variants of the Latent Diffusion Model (LDM) [37], specifically Stable Diffusion models. LDMs employ cross-attention layers to condition diffusion models with text embeddings, necessitating a mapping of target subject images to latent spaces compatible with the diffusion models, either as text embeddings or as supplementary model parameters. This process involves backpropagation through the entire diffusion model, often comprising over a billion parameters, contributing to the inefficiency of existing P-T2I methods.
In contrast, the UnCLIP [35] T2I family of LDMs offers an intriguing alternative, wherein text embeddings are mapped to vision embeddings through a diffusion transformer prior network before passing the conditioning to the diffusion UNet image generator. This mechanism, as illustrated in Figure 2, enables the generator to accurately reproduce original images from their visual embeddings alone. Consequently, our research pivots towards leveraging the UnCLIP T2I stack for P-T2I strategies. However, like their predecessors, these T2I diffusion priors are also parameter-heavy, with around 1 billion parameters. Until recently, ECLIPSE [31] showed that it is possible to compress this 1B parameter of the T2I prior model down to 33M by removing the diffusion modeling.
Table 1: A quick overview of previous works on T2I personalization. Our method is the first to offer multi-concept personalization without depending on diffusion UNet models (except for inference). We provide the extended overview table in the supplementary materials.
Method Multi Finetuning Diffusion Total opt. GPU
Concepts Free Free params Hours
Textual Inversion [12] 768 1
DreamBooth [38] 0.9B 0.5
ELITE [48] 77M -
BLIP-Diffusion [21] 1.5B 2304
IP-Adapter [50] 22M -
[…]
Subject-Diffusion [27] 252M -
Kosmos-G [29] 1.9B 12300
$\lambda$-ECLIPSE (ours) 34M 74
Figure 2: CLIP(vision) features capture the semantics and fine-grained visual details. UnCLIP models (e.g., Kandinsky v2.2) can effectively leverage this latent space to reconstruct the image from its embedding.
ECLIPSE posits that text-to-image mapping can be optimized through contrastive pre-training. It inputs text embeddings and estimates corresponding image embeddings, ensuring strong alignment with textual features. Building on these insights, to enhance this framework and deepen the comprehension of novel visual subjects, we introduce a novel pre-training strategy and propose the $\lambda$-ECLIPSE  T2I prior model for the UnCLIP framework. Specifically, we propose a subject-driven instruction tuning task involving the image-text interleaved data. This involves creating 1.6 million high-quality image-text pairs, where text tokens linked to subjects are substituted with image embeddings. Subsequently, $\lambda$-ECLIPSE  is trained to estimate image embeddings that not only harmonize with text semantics but also encapsulate subject representations.
This model facilitates multi-subject-driven image generation using UnCLIP models without imposing excessive resource demands. $\lambda$-ECLIPSE  bypasses the reliance on diffusion latent space and operates within a pre-trained CLIP latent space [33]. We leverage Kandinsky v2.2 diffusion UNet model’s [36] capacity to interpret image structures from CLIP embeddings (refer to Figure 2). Importantly, CLIP image embeddings even accurately capture the details of the canny edge map [6]. Therefore, we extend $\lambda$-ECLIPSE to incorporate canny edge maps as auxiliary guides for subject-driven T2I. Our extensive experimental validation demonstrates that $\lambda$-ECLIPSE  achieves unparalleled performance in single and multi-subject composition alignment while preserving the integrity of reference subjects, competitively rivaling Multimodal Large Language Models (MLLM) based methods that require intense computing resources. At last, we find that $\lambda$-ECLIPSE inherits the continuous latent space from the CLIP; for the first time allowing the interpolations between multiple concepts for P-T2I models. Figure 1 illustrates the qualitative capabilities of $\lambda$-ECLIPSE.
In summary, our contributions are: (1) we introduce a straightforward and yet very effective diffusion-free multi-subject personalization strategy by estimating subject-aligned image embeddings in CLIP(vision) latent space, (2) despite significantly reducing resource requirements, our method achieves state-of-the-art performance, and (3) we demonstrate the potential of utilizing diffusion image generators as standalone pre-trained entities, enabling controlled outputs without necessitating explicit fine-tuning.
2 Related Works
Figure 3: This figure illustrates the three stages of the $\lambda$-ECLIPSE pipeline. 1) Create the image-text interleaved features by utilizing the pre-trained CLIP encoders. 2) Pre-train the $\lambda$-ECLIPSE (34M parameters) using Eq. 1, which ensures the mapping to the desired latent space given the image-text interleaved data. During inference, we just need a single-forward pass to estimate the corresponding image embedding. 3) During inference, the Kandinsky v2.2 diffusion UNet model takes the output from the $\lambda$-ECLIPSE (prior) and generates the corresponding images.
Text-to-Image Generative Models.
Pioneering efforts in image generation, notably DALL-E [34] and CogView [11], leveraged autoregressive models to achieve significant results. Recent advancements predominantly feature diffusion models, acclaimed for their high image fidelity and diversity in text-to-image (T2I) generation. A notable example is Stable Diffusion, which builds upon the Latent Diffusion Model (LDM) [37] and excels in semantic and conceptual understanding by transitioning training to latent space. Imagen [40], Pixart-$\alpha$ [8], and DALL-E 3 [5] propose using a large T5 language model to improve language understanding. DALL-E 2 [35] along with its UnCLIP variation models such as Kandinsky [36] and Karlo, uses a diffusion prior and diffusion UNet modules to generate images using the pre-trained CLIP [33] model.
Personalized T2I Methods.
Approaches like Textual Inversion [12], DreamBooth [38], and Custom Diffusion [20] focus on training specific parameters to encapsulate visual concepts. LoRA [16] and Perfusion [45] target efficient fine-tuning adjustments, particularly rank 1 modifications. However, these methods are constrained by their requirement for concept-specific tuning. ELITE [48] was the first approach addressing fast personalization for single-subject T2I. BLIP-Diffusion [21] adapts the BLIP2 encoder [22], training approximately 1.5B parameters to enable zero-shot, subject-driven image generation. IP-Adapter [50] introduces a decoupled cross-attention mechanism, negating the need to train the foundational UNet model by permitting fine-tuning of a reduced parameter set (22M).
Mix-of-Show [14] and Zip-LoRA [42] train individual concepts and then combine them to generate multiple subjects. Break-A-Scene [4] shows multi-concept capability but requires single images containing diverse objects. Subject Diffusion [27] creates a high-quality dataset and presents the precision control for fast multi-subject image generation. Kosmos-G [29], akin to Subject-Diffusion, employs a Multimodal Large Language Model (MLLM) for text-image embedding alignment, though it necessitates extensive parameter optimization (1.9B). These multi-subject P-T2I methods are not only demanding in terms of parameters but also depend on a massive number of frozen parameters of the diffusion UNet model, increasing training computational loads. In contrast, our model, $\lambda$-ECLIPSE, forgoes test-time fine-tuning and training-time reliance on the diffusion UNet model for single and multi-concept, control-guided P-T2I, positioning it as a resource-efficient solution.
At last, methods like GLIGEN [23], ControlNet [51], and UniControl [32] incorporate additional controls (i.e., edge map, depth, segmentations) into the diffusion model to generate the desired images. BLIP-Diffusion, IP-Adapter, and Kosmos-G can leverage such pre-trained controls. However, in many scenarios, these controls are too strong which makes generated images lose subject-specific details. While $\lambda$-ECLIPSE can be plugged with these pre-trained methods, we explore the possibility of learning canny edge as the additional guidance for P-T2I. We show that $\lambda$-ECLIPSE learns to balance the edge map, subjects, and text composition.
3 Method
In this section, we introduce $\lambda$-ECLIPSE, our approach to multi-subject text-to-image personalization. This method extends the capabilities of the existing ECLIPSE  framework through a novel image-text interleaved pretraining strategy, notably omitting the need for explicit diffusion modeling. Our approach mainly capitalizes on the efficient utilization of the CLIP latent space. Figure 3 outlines the end-to-end framework.
The primary objective of $\lambda$-ECLIPSE is to facilitate both single and multi-subject T2I generation processes, accommodating controlled conditional guidance such as edge maps. Initially, we detail the problem formulation and elaborate on the extended design of the ECLIPSE  T2I prior module. Subsequently, we delve into the image-text interleaved training methodology. This fine-tuning process enables the ECLIPSE  model to harness semantic correlations between CLIP image and text latent spaces while preserving subject-specific visual features. Upon attaining the target embedding, the diffusion UNet model, derived from Kandinsky v2.2, is employed to reconstruct the image.
3.1 Text-to-Image Prior Mapping
In the UnCLIP T2I models, the objective of the text-to-image prior model ($f_{\theta}$) is to establish a proficient text-to-image embedding mapping. This model is designed to adeptly map textual representations to their corresponding visual embeddings, denoted as ($f_{\theta}:z_{y}\rightarrow z_{x}$), where $z_{x/y}$ represent the embeddings for images and text, respectively. The visual embedding predictions ($\hat{z_{x}}=f_{\theta}(z_{y})$) are then effectively utilized by the diffusion image generators ($h_{\phi}$), which are inherently conditioned on these vision embeddings ($h_{\phi}:z_{x}\rightarrow x$).
As illustrated in Figure 2, these diffusion image generators possess a remarkable capability for precise image reproduction, since $z_{x}$ encompasses comprehensive image information. Our goal is to accurately estimate the image embedding $\hat{z}_{x}$, incorporating the subject representations, thereby eliminating reliance on $h_{\phi}$ during training. Existing LDM-based P-T2I methods are limited by the LDM’s singular module approach ($h_{\phi}:z_{y}\rightarrow x$). Consequently, mastering the latent space of $h_{\phi}$ becomes essential for effective P-T2I for the baseline methodologies.
We propose a mapping function, $f^{\prime}_{\theta}$, which processes text representations ($z_{y}$) alongside subject ($x_{i}$) specific representations ($z_{x_{i}}$), to derive an image embedding that encapsulates both text prompts and subject visuals ($\hat{z}_{x}$). The challenge lies in harmonizing $z_{x_{i}}$ and $z_{y}$ within $f^{\prime}_{\theta}:(z_{y},z_{x_{i}})\rightarrow\hat{z}_{x}$, ensuring alignment while preventing overemphasis on either aspect, as this could compromise composition alignment. To address this, we employ the contrastive pre-training strategy after [31]:
$\begin{split}\mathcal{L}_{prior}=&\underset{\begin{subarray}{c}\epsilon\sim\mathcal{N}(0,I)\\ z_{y},z_{x}\end{subarray}}{\operatorname{\mathbb{E}}}\Big{[}||z_{x}-f_{\theta}(\epsilon,z_{y})||_{2}^{2}\Big{]}\\ &-\frac{\lambda}{N}\sum_{i=0}^{N}\log\frac{\exp(\langle\hat{z}^{i}_{x},z^{i}_{y}\rangle/\tau)}{\sum_{j\in[N]}\exp(\langle\hat{z}^{i}_{x},z^{j}_{y}\rangle/\tau)}.\end{split}$ (1)
Here, $\lambda$ serves as the hyperparameter. The first loss term measures the mean-squared error between the estimated and actual image embeddings, primarily ensuring concept alignment. However, our preliminary studies reveal that exclusive reliance on this term diminishes composition alignment, echoing observations made by [31]. Therefore, we stick with the contrastive loss component to bolster compositional generalization, with $\lambda$ balancing concept and composition alignment.
Figure 4: Qualitative examples of $\lambda$-ECLIPSE.
Figure 5: This figure illustrates the qualitative comparison of $\lambda$-ECLIPSE w.r.t. the state-of-the-art approaches for single-subject T2I generations. We choose concepts and prompts from the Dreambench dataset. For all methods, concepts, and prompts, we generate the four images and select the best one that represents the concept and composition more accurately.
Additional Control-based T2I Prior Mapping.
Acknowledging the limitations in existing methods, which necessitate learning the diffusion latent space even for additional control inputs, we endeavor to achieve a more nuanced balance between subject, text, and supplementary conditions. Consequently, we have augmented $\lambda$-ECLIPSE to accommodate an additional modality, a canny edge map, providing more refined control over subject-driven image generation. This entails modifying the prior model to accept additional conditions ($f_{\theta}^{\prime}:(z_{y},z_{x_{i}},z_{c})\rightarrow\hat{z}_{x}$, where $z_{c}$ symbolizes the additional modality embedding). During training, we drop $z_{c}$ for 1% to improve the unconditional generations.
Incorporating canny edge maps during training enhances stability and broadens the generalization capabilities of $\lambda$-ECLIPSE, yielding benefits even in the absence of these controls during inference. Our findings indicate that $\lambda$-ECLIPSE successfully learns a unified mapping function, effectively estimating target image representations from a synergy of text, subjects, and edge map.
3.2 Image-text Interleaved Training
Our approach targets developing a versatile prior model capable of processing diverse inputs to estimate visual outputs. Drawing from earlier methodologies, a straightforward solution involves concatenating different inputs, like combining text (“a dog wearing sunglasses”) with respective concept-specific images. However, preliminary experiments indicated that this method does not effectively capture the intricate relationships between target text tokens (e.g., “dog”) and the corresponding concept images.
To address this, we adopt the interleaved pre-training strategy used in Kosmos-G, but with a notable modification to enhance resource efficiency. We incorporate pretrained CLIP text and vision encoders for extracting modality-specific embeddings—separating text-only from subject-specific images. The key refinement in our process is the substitution of subject token embeddings with corresponding vision embeddings instead of introducing additional tokens to handle the image embeddings via resampler [2]. This alteration allows us to bypass the need to train the prior or MLLMs, significantly improving the model’s proficiency in handling interleaved data.
For the generation of high-quality training datasets, we carefully selected 2 million high-quality images from the LAION dataset [41], each with a resolution of 1024x1024. Utilizing BLIP2, we generate captions for these images and employ SAM [19] for extracting noun or subject-specific segmentation masks. Given the CLIP model’s requirement for 224x224 resolution images, we avoid resizing the masks within their original resolutions. Instead, we opt for cropping the area of interest using DINO [24], followed by resizing the masked object while preserving its aspect ratio. This technique is crucial in retaining maximum visual information for each subject during the training phase. We provide more details about the filters used in the supplementary material.
In summary, the ECLIPSE  prior, trained with our image-text interleaved data and supplementary conditions, presents an efficient pathway for P-T2I. The culmination of this process yields the $\lambda$-ECLIPSE model, a refined version characterized by enhanced effectiveness in handling and interpreting multimodal data.
4 Experiments
In this section, we first introduce the experimental setup, evaluations, and baselines. Later, we delve into the qualitative and quantitative results.
Training and inference details.
We initialize our model, $\lambda$-ECLIPSE, equipped with 34M parameters. We train our model on an image-text interleaved dataset of 2M instances, partitioned into 1.6M for training and 0.4M for validation. The model is specifically tuned for the Kandinsky v2.2 diffusion image decoder. Therefore, we use pre-trained OpenCLIP-ViT-G/14222https://huggingface.co/laion/CLIP-ViT-g-14-laion2B-s12B-b42K as the text and vision encoders, ensuring alignment with Kandinsky v2.2 image embeddings. Training is executed on 2 x A100 GPUs, leveraging a per-GPU batch size of 512 and a peak learning rate of 0.00005, across approximately 100,000 iterations, summing up to 74 GPU hours. During inference, the model employs 50 DDIM steps and 7.5 classifier-free guidance for the Kandinsky v2.2 diffusion image generator. Adhering to baseline methodologies, we perform the P-T2I following the baseline papers’ protocols. For $\lambda$-ECLIPSE, target subject pixel regions in reference images are segmented before embedding extraction via the CLIP(vision) encoder. We drop the canny edge map during inference unless specified explicitly.
Evaluation setup.
We primarily utilize Dreambench (encompassing 30 unique concepts with 25 prompts per concept) for qualitative and quantitative evaluations using DINO and CLIP-based metrics [38]. Due to their limitations, we extend our evaluations on the ConceptBed [30] benchmark (covering 80 diverse imagenet concepts and a total of 33K composite prompts), where we report performance on concept replication, concept, and composition alignment using the Concept Confidence Deviation ($CCD(\downarrow)$) metric.
4.1 Results & Analysis
Table 2: Quantitative comparisons of different methodologies on Dreambench. The Bold and underline represent the metric-specific first and second-ranked methods, respectively. * represents that we reproduce the open-source methodologies. Other results are borrowed from the Subject-Diffusion [27].
Method Base Model DINO (normal-$\uparrow$) CLIP-I (normal-$\uparrow$) CLIP-T (normal-$\uparrow$)
Textual Inversion SDv1.5 0.569 0.780 0.255
DreamBooth SDv1.5 0.668 0.803 0.305
Custom Diffusion SDv1.5 0.643 0.790 0.305
Re Imagen Imagen 0.600 0.740 0.270
ELITE SDv1.4 0.621 0.771 0.293
Subject-Diffusion SDv1.5 0.711 0.787 0.293
BLIP-Diffusion* SDv1.5 0.603 0.793 0.291
IP-Adapter* SDv1.5 0.629 0.827 0.264
IP-Adapter* SDXL 0.613 0.810 0.292
Kosmos-G* SDv1.5 0.618 0.822 0.250
[…]
Figure 6: Multi-subject qualitative examples illustrating the comparison between $\lambda$-ECLIPSE and other SOTA baselines.
Table 3: Quantitative comparisons of different methodologies on ConceptBed. We present results on $CCD~{}(\downarrow)$ across three evaluation categories. The Bold and underline represent the metric-specific first and second-ranked methods, respectively. * represents that we reproduce the open-source methodologies. Other results are borrowed from the ConceptBed [30].
Method Base Concept Concept Composition
Model Replication Alignment Alignment
Textual Inversion SDv1.4 0.0662 0.1163 0.1436
Dreambooth SDv1.4 0.0880 0.3551 0.0360
Custom Diffusion SDv1.4 0.2309 0.4882 0.0204
ELITE* SDv1.4 0.3195 0.4666 0.1832
BLIP-DIffusion* SDv1.5 0.3510 0.3245 0.1589
IP-Adapter* SDXL 0.3665 0.3571 0.0641
[…]
Figure 7: Qualitative examples for edge-guided P-T2I showcasing the comparison between BLIP-Diffusion vs. $\lambda$-ECLIPSE.
Quantitative comparison.
The quantitative assessments are detailed in Table 2 and Table 3, focusing on the single-concept T2I task. For Dreambench, we generate and evaluate four images per prompt, reporting average performance on three metrics (DINO, CLIP-I, and CLIP-T). In the case of ConceptBed, we process each of the 33K prompts to generate a single image. The results, as depicted in these tables, highlight $\lambda$-ECLIPSE’s superior performance in composition tasks and its competitive edge over even fine-tuning-based methodologies, while maintaining robust concept alignment. $\lambda$-ECLIPSE exhibits a notable proficiency in concept replication, albeit with a marginal trade-off in concept alignment for enhanced composition fidelity. Comparatively, IP-Adapter, ELITE, and BLIP-Diffusion prioritize concept alignment, often at the expense of composition alignment. Noteworthy is $\lambda$-ECLIPSE’s efficiency, achieved with significantly fewer resources.
Qualitative comparisons.
In Figure 5, we present a range of single subject-specific images generated by various methodologies including BLIP-Diffusion, IP-Adapter, Kosmos-G, and $\lambda$-ECLIPSE. $\lambda$-ECLIPSE demonstrates exemplary proficiency in composition while ensuring concept alignment. In contrast, the baselines often overemphasize reference images (notably in rows 2 and 4) or exhibit concept dilution (as seen in rows 6 and 8), leading to higher concept alignment but compromised composition. Additionally, Figure 6 exhibits $\lambda$-ECLIPSE’s multi-concept generation prowess, in comparison to ZipLoRA (fine-tuning-based approach) and Kosmos-G (Multimodal LLM-based approach), underscoring its capability to rival compute-intensive methodologies.
Figure 2 previously demonstrated Kandinsky v2.2’s ability to partially reconstruct canny edge maps while preserving semantic integrity. Building on this, Figure 7 offers qualitative insights, comparing BLIP-Diffusion and $\lambda$-ECLIPSE in edge-guided subject-driven T2I scenarios. Here, BLIP-Diffusion adheres strictly to the imposed conditions, often at the cost of concept retention (rows 1, 3, and 4). Conversely, $\lambda$-ECLIPSE not only complies with these conditions but also adeptly balances the need for high concept and composition alignment.
Figure 8: Interpolation between four concepts. Here, we estimate the image embedding using $\lambda$-ECLIPSE corresponding to each corner and then interpolate from top to bottom and left to right. At last, we use the Kandinsky v2.2 diffusion UNet model to generate the images with fixed random seeds from these sets of image embeddings.
Table 4: Ablation studies w.r.t. to the key components of $\lambda$-ECLIPSE  design. We report the concept and composition alignment for single-subject T2I using $CCD~{}(\downarrow)$ on the ConceptBed benchmark.
Model Concept Composition
Alignment Alignment
Projection loss 0.394 0.008
w/ contrastive loss (lambda=0.5) 0.435 -0.043
w/ contrastive loss (lambda=0.2) 0.402 -0.026
[…]
Ablations.
We extend our study to evaluate the individual contributions of different components in $\lambda$-ECLIPSE. Initially, the model’s performance with solely the projection loss (referenced in Eq.1) is assessed. Subsequent experiments involve training $\lambda$-ECLIPSE variants with varying hyperparameters for the contrastive loss, specifically $\lambda$ values of 0.2 and 0.5. A comparative analysis of these baselines is conducted against the fully equipped $\lambda$-ECLIPSE model, which incorporates $\mathcal{L}_{prior}$ (Eq.1) with $\lambda=0.2$ and utilizes canny edge maps during training. As Table 4 illustrates, relying solely on projection loss results in high concept alignment but compromises compositions. The contrastive loss variant with $\lambda=0.5$ enhances composition alignment at the expense of concept alignment, whereas $\lambda=0.2$ achieves a more balanced performance. Crucially, the integration of canny edge maps during training optimally balances both alignments and, specifically, improves the concept alignment. Here, the negative values indicate that the $CCD$ oracle model performs similarly to real images having the same set of compositions.
Multi-subject interpolation.
A key attribute of the CLIP latent space is the ability to perform smooth interpolation between two sets of embeddings. We conducted experiments to demonstrate $\lambda$-ECLIPSE’s ability to learn and replicate this smooth latent space transition. We selected two distinct dog breeds ($<$dog1$>$, $<$dog2$>$) and two types of hats ($<$hat1$>$, $<$hat2$>$) as the concepts. $\lambda$-ECLIPSE was then used to estimate image embeddings for all four possible combinations, each corresponding to the input phrase “a $<$dog$>$ wearing a $<$hat$>$.” Figure 8 showcases a gradual and seamless transition in the synthesized images from the top left to the bottom right. Unlike current diffusion models, which often exhibit sensitivity to input variations necessitating numerous iterations for desired outcomes, $\lambda$-ECLIPSE inherits CLIP’s smooth latent space. This inheritance not only facilitates progressive changes in the conceptual domain but also extends the model’s utility by enabling personalized multi-subject interpolations, a feature that is not inherently available in previous P-T2I methods.
5 Conclusion
In this paper, we have introduced a novel diffusion-free methodology for personalized text-to-image (T2I) applications, utilizing the latent space of the pre-trained CLIP model with high efficiency. Our $\lambda$-ECLIPSE model, trained on an image-text interleaved dataset of superior quality, marks a significant advancement in the field. This model outperforms existing methods by its capability to execute single-concept, multi-concept, and edge-guided controlled P-T2I tasks using a singular model framework, while simultaneously minimizing resource utilization. Notably, $\lambda$-ECLIPSE sets a new benchmark in achieving state-of-the-art results in composition alignment, coupled with exceptional maintenance of subject fidelity. Furthermore, our research illuminates the potential of $\lambda$-ECLIPSE in exploring and leveraging the smooth latent space. This capability unlocks new avenues for interpolating between multiple concepts and their amalgamation, thereby generating entirely novel concepts. Our findings underscore a promising pathway where pre-trained diffusion image generators can be effectively controlled without necessitating extra supervision.
Acknowledgments
This work was supported by NSF RI grants #1750082, #2132724, and CPS grant #2038666. The views and opinions of the authors expressed herein do not necessarily state or reflect those of the funding agencies and employers.
Appendix A Preliminaries for T2I Diffusion Models
As evidenced in numerous contemporary studies regarding T2I models, Stable Diffusion (SD) has emerged as a predominant backbone for text-to-image (T2I) models. SD involves training diffusion models in latent space, reversing a forward diffusion process that introduces noise into the image. A notable feature of SD is its integration of cross-attention, facilitating various conditions like text input. Operating in VQVAE latent space, SD not only achieves exceptional generative performance surpassing that in pixel space but also significantly reduces computational demands.
UnCLIP models (such as DALL-E 2) are very similar to the Stable Diffusion. In contrast, the UnCLIP stack takes the modular approach. UnCLIP first trains the diffusion text-to-image to the image prior ($f_{\theta}$) to estimate the image embeddings ($z_{x}$) from the text embeddings ($z_{y}$). In parallel, a UNet-like diffusion image generator ($h_{\phi}$) is trained to generate images ($x$) conditioned on ground truth vision embeddings ($z_{x}$).
Traditionally, T2I prior is modeled to estimate $x_{0}$-prediction instead of $\epsilon$-prediction. Given the forward function $z_{x}^{(t)}\sim q(t,z_{x})$, the goal of $f_{\theta}$ is to directly estimate $z_{x}$ for all timesteps $t\sim[0,T]$ as:
$\mathcal{L}_{prior}=\underset{\begin{subarray}{c}t\sim[0,T],\\ z_{x}^{(t)}\sim q(t,z_{x})\end{subarray}}{\operatorname{\mathbb{E}}}\Big{[}||z_{x}-f_{\theta}(z_{x}^{(t)},t,z_{y})||_{2}^{2}\Big{]}.$ (2)
ECLIPSE presents the contrastive learning strategy (Eq. 1) instead of minimizing Eq. 2. The diffusion image generator is trained by following standard $\epsilon$-prediction formulation. Here, $h_{\phi}$ will estimate the ground truth added Gaussian noise $\epsilon\sim N(0,I)$, given the noise image $X^{(t)}$ for all timesteps $t\sim[0,T]$ and input conditions (such as $z_{x}$, $z_{y}$).
$\mathcal{L}_{decoder}=\underset{\begin{subarray}{c}\epsilon\sim N(0,I)\\ t\sim[0,T],\\ (z_{x},~{}z_{y})\end{subarray}}{\operatorname{\mathbb{E}}}\Big{[}||\epsilon-h_{\phi}(x^{(t)},t,z_{x},z_{y})||_{2}^{2}\Big{]}.$ (3)
For models like Kandinsky v2.2, we drop the $z_{y}$ to condition the model on $z_{x}$. Therefore, $\lambda$-ECLIPSE also only conditions the image generation with $z_{y}$ in the prior stage.
Appendix B Image-Text Interleaved Training Details
Dataset Creation
In constructing the dataset, we adhered to the data processing pipeline of Subject Diffusion [27]. We utilized the LAION-5B High-Res dataset, requiring a minimum image size of 1024x1024 resolution. Original captions were replaced with new captions generated by BLIP-2 (flan-t5-xl) [^3] https://huggingface.co/Salesforce/blip2-flan-t5-xl. Subjects were extracted using Spacy [^4] https://spacy.io. For each subject, we identified bounding boxes employing Grounding DINO [24], setting both box-threshold and text-threshold values to 0.2. We retained images with 1 to 8 detected bounding boxes, discarding the rest. Additionally, captions with multiple instances of identical objects were filtered, allowing a maximum of 6 identical objects. Following bounding box detection, individual subject masks were isolated using Segment-Anything (SAM) [19]. To enhance the efficiency of the training process, we pre-processed the dataset by pre-extracting features from CLIP vision and text encoders. During this phase, images predominantly featuring a background (white portion) exceeding 10% of the total area were excluded. We preserved bounding boxes with a width-height ratio ranging from 0.08 to 0.7 and logit scores of at least 0.3. Masks constituting less than 40% of the bounding box area were discarded. For the selection of subjects in images, we constrained the range to 1-4 subjects per image, excluding those with more than 4 subjects. At last, the interleaved image-text examples with respective ground truth images are shown in Figure 9.
Figure 9: Examples of image-text interleaved training data. Left column shows the input the prior model and right images shows the ground truth corresponding images. Note: these examples are generated from $\lambda$-ECLIPSE for better interpretability.
Dataset Statistics
In the final analysis, our dataset comprised a total of 1,990,123 images. The distribution of subjects per image exhibited a range from 1 to 4, with the following breakdown: 1,479,785 images featuring one subject, 432,831 images with two subjects, 65,597 images containing three subjects, and 11,910 images showcasing four subjects. The overall count of unique subjects acquired from this dataset amounted to 30,358. We partitioned our dataset into an 80:20 split between training and validation, reserving the remaining 1.6 million images for training and the rest for validation.
Table 5: The detailed overview of subject-driven text-to-image generative methodologies. * represents the backbone base models listed are subject to potential updates or modifications.
Method Multi-Concept Finetuning Free Base Model Image Input
Re-Imagen [9] Imagen Single
Textual Inversion [12] SDv1.4 Multiple
DreamBooth [38] SDv1.4 Multiple
Custom Diffusion [20] SDv1.4 Multiple
ELITE [48] SDv1.4 Single
E4T [13] SD Single
Cones [25] SDv1.4 Single
SVDiff [15] SD Multiple
UMM-Diffusion [28] SDv1.5 Single
XTI [47] SDv1.4 Multiple
Continual Diffusion [44] - Multiple
InstantBooth [43] SDv1.4 Multiple
SuTi [10] Imagen Multiple
Taming [17] Imagen Single
BLIP-Diffusion [21] SDv1.5 Single
Cones 2 [26] SDv2.1 Single
DisenBooth [7] SDv2.1 Single
FastComposer [49] SDv1.5 Single
Perfusion [45] SDv1.5 Multiple
Mix-of-Show [14] Chilloutmix Multiple
NeTI [1] SDv1.4 Mulitple
Break-A-Scene [4] SDv2.1 Single*
ViCo [46] SDv1.4 Mulitple
Domain-Agnostic [3] - Single
Subject-Diffusion [27] SDv2 Single
HyperDreamBooth [39] SDv1.5 Single
IP-Adapter [50] SDv1.5 Single
Kosmos-G [29] SDv1.5 Single
Zip-LoRA [42] sdxl Multiple
CatVersion [53] SDv1.5 Multiple
SSR-Encoder [52] SDv1.5 Single
[…]
Appendix C Implementation Details
The $\lambda$-ECLIPSE transformer prior architecture is significantly more compact compared to other Text-to-Image (T2I) methodologies. Our model employs a configuration of 16 Attention Heads, each with a dimension size of 32, alongside a total of 10 layers. Additionally, the embedding dimension size for our model is set at 1280, supplemented by 4 auxiliary embeddings (including, one for canny edge map). As $\lambda$-ECLIPSE is not a diffusion prior model, we do not keep time embedding layers. Overall, the $\lambda$-ECLIPSE model comprises approximately 34 million parameters, establishing it as a streamlined yet effective solution for Personalaized-T2I. Notably, the standard UnCLIP T2I priors contain 1 billion parameters.
Appendix D Extended P-T2I Baselines Comparison
We further expand our comparative analysis of P-T2I methods encompassing a total of 32 approaches including ours and parallel works. Table 5 summarizes them into four crucial aspects: 1) multi-concept support, 2) fine-tuning free, 3) base model types, and 4) the required number of input images. To summarize, $\lambda$-ECLIPSE is the only methodology built on top of the UnCLIP models while supporting multi-concept personalization, fine-tuning free, and only requires a single reference concept image. We detail the comparison below:
Multi-Concept Generation.
Multi-concept generation enables users to integrate multiple personal subjects to generate an image that follows the text prompts and aligns with all the concept visuals. Of the 32 methods analyzed, 14 offer this capability. Ten methods require separate training for each concept and then an additional fusing step for combining the learned concepts. (Mix-of-Show). While only four methods support fast multi-concept personalization. Only a few can learn auxiliary guided information such as canny edge, depth maps, or open-pose and adapt style variation.
Fine-tuning Free (Fast Personalization).
Many models require test-time fine-tuning. Each varies on which part alteration occurs, as early models tend to modify the whole UNet. In contrast, recent models tune a small portion of the cross-attention layers or introduce additional layers performing as adapters. In our analysis of P-T2I models, 13 out of 32 models employ a fine-tuning free approach which enables fast personalization.
Diffusion Independent.
A majority of the reviewed models utilize diffusion models, with Stable Diffusion being the predominant choice, spanning versions 1.4, 1.5, 2.1, and XL. Few adapt Imagen (SuTi, Taming) and interestingly Mix-of-show employs ChillOutMix as their pre-trained model, known for its adeptness at preserving realistic concepts like human faces. A unique outlier in this landscape is our $\lambda$-ECLIPSE, the only one that eschews the use of any diffusion prior model.
Easiness to Use (Input and Output).
Ease of use, particularly regarding model input and output, significantly impacts user convenience in subject-driven generalization. A more user-friendly model typically requires a single input image, as opposed to multiple images of the same subject. In our study, 18 models offer P-T2I capabilities with just one input image. In contrast, others often require 4 to 5 images of the subject. Additionally, some models necessitate storage space for learned concepts, ranging from a few hundred kilobytes (e.g., Perfusion, HyperDreamBooth) to several megabytes (e.g., Zip-LoRA). Our model stands out by eliminating the need for individual concept pre-learning or storing any artifacts for P-T2I utilization, offering a streamlined, efficient user experience.
Appendix E More Qualitative Results
In this section, we present an array of extended examples of P-T2I generation, which presents complex compositions that are hard to achieve in $\lambda$-ECLIPSE as well as other models. Figure 10 showcases the complexity level that notably escalates as it goes from top to bottom. As the complexity of the visual concepts increases, we observe that all methodologies perform poorly in maintaining the subject-fidelity. This applies to $\lambda$-ECLIPSE as well. Interestingly, $\lambda$-ECLIPSE retains the compositional understanding while baselines perform poorly in all scenarios.
In addition, we demonstrate some examples of how P-T2I methods exhibit inconsistency in their generated results depending on each trial. Figure 11 exhibits consistency in single and multiple concept generation across both models. However, Kosmos-G exhibits inconsistency in generating multiple concepts, sometimes placing the Ironman suit in various parts of the dog or omitting it entirely. This also means that $\lambda$-ECLIPSE reduces the diversity of the generated images to improve the consistency. However, this behavior is observed across the UnCLIP family of models.
In our final exploration, we examine the preservation of actual human facial characteristics in target scenarios, combined with varying captions. While each model attempts to retain the original face, none can precisely replicate the exact personal facial features. These examples generally fail to accurately represent the intended compositions, except in one instance in IP-Adapter FaceID.
Appendix F Limitations
Our work represents a pioneering effort in harnessing a method that utilizes the latent space of pre-trained CLIP models for P-T2I generation, it is important to acknowledge certain limitations. Firstly, CLIP, despite its capabilities, does not perfectly capture hierarchical representations, and this limitation can occasionally result in suboptimal outcomes. Consequently, the CLIP contrastive loss can cause the model to deviate from the original subject features, making performing P-T2I on complex concepts, such as human faces, more challenging. We anticipate that the enhancement of CLIP representations will correspondingly improve the performance of our framework in the realm of P-T2I mapping. The $\lambda$-ECLIPSE model has been trained with 34 million parameters and 1.6 million images. However, increasing the data quality and parameters might lead to improvements in the results.
Appendix G Broader Impact
The recent retraction of the LAION-5B dataset, prompted by concerns over the presence of data samples potentially breaching ethical guidelines, has significantly affected various research initiatives. This situation underscores the crucial importance of careful dataset selection and raises a cautionary note for users about the ethical dimensions involved. In line with this, the community needs caution in the application of associated methodologies as well. Subject-driven image generation or Personalized Text-to-Image (P-T2I) methods have the potential to be a transformative tool in numerous domains.
For their positive influence, they enable users to effortlessly generate, modify, and synthesize original subjects into diverse environments, thereby enriching creative expression. On the other hand, the ease of altering and creating images raises concerns about the responsible use of this technology which requires significant ethical and legal considerations. We recommend developers provide a more secure way (such as image attribution) for end-users to ensure accountability for misuse of such models [18]. Users must be acutely aware of being able to infringe intellectual property rights and create misleading or harmful content. As such, those employing subject-driven image-generation techniques should exercise careful judgment, ensuring that their work adheres to ethical standards and legal boundaries. It is imperative that the broader implications of this technology are considered, and that a commitment to responsible and conscientious use guides its application.
Figure 10: Qualitative examples of increasing complexity of novel visual concepts as we move from top to bottom.
Figure 11: Qualitative examples of showcasing the consistency comparisons between Kosmos-G and $\lambda$-ECLIPSE.
Figure 12: Qualitative examples of showcasing the failure cases on human faces on Kosmos-G, IP-Adapter (SDXL), IP-Adapter (FaceID), and $\lambda$-ECLIPSE.

### 09 Apr 2024 v2

@@
-by Leveraging CLIP Latent Space
+Maitreya $*{}^{\dagger}$   Sangmin Jung*   Chitta Baral   Yezhou Yang
 Abstract
-* indicates equal contribution, $\dagger$ Corresponding Author: maitreya.patel@asu.edu.
-Despite the recent advances in personalized text-to-image (P-T2I) generative models, subject-driven T2I remains challenging.
-The primary bottlenecks include 1) Intensive training resource requirements, 2) Hyper-parameter sensitivity leading to inconsistent outputs, and 3) Balancing the intricacies of novel visual concept and composition alignment.
-We start by re-iterating the core philosophy of T2I diffusion models to address the above limitations.
-Predominantly, contemporary subject-driven T2I approaches hinge on Latent Diffusion Models (LDMs), which facilitate T2I mapping through cross-attention layers.
+Despite the recent advances in personalized text-to-image (P-T2I) generative models, it remains challenging to perform finetuning-free multi-subject-driven T2I in a resource-efficient manner.
+Predominantly, contemporary approaches, involving the training of Hypernetworks and Multimodal Large Language Models (MLLMs), require heavy computing resources that range from 600 to 12300 GPU hours of training.
+These subject-driven T2I methods hinge on Latent Diffusion Models (LDMs), which facilitate T2I mapping through cross-attention layers.
 While LDMs offer distinct advantages, P-T2I methods’ reliance on the latent space of these diffusion models significantly escalates resource demands, leading to inconsistent results and necessitating numerous iterations for a single desired image.
-Recently, ECLIPSE has demonstrated a more resource-efficient pathway for training UnCLIP-based T2I models, circumventing the need for diffusion text-to-image priors.
-Building on this, we introduce $\lambda$-ECLIPSE [^1] The designation $\lambda$-ECLIPSE is inspired by its conceptual alignment with the $\lambda$-calculus.
-In this context, the $\lambda$-ECLIPSE model functions similarly to a functional abstraction within $\lambda$-calculus, where it effectively binds variables.
-These variables, in our case, represent novel visual concepts that are integrated through composition prompts..
-Our method illustrates that effective P-T2I does not necessarily depend on the latent space of diffusion models.
-$\lambda$-ECLIPSE  achieves single, multi-subject, and edge-guided T2I personalization with just 34M parameters and is trained on a mere 74 GPU hours using 1.6M image-text interleaved data.
-Through extensive experiments, we also establish that $\lambda$-ECLIPSE  surpasses existing baselines in composition alignment while preserving concept alignment performance, even with significantly lower resource utilization.
-Figure 1: Our $\lambda$-ECLIPSE  can estimate subject-specific image embeddings while maintaining the balance between concept and composition alignment.
-$\lambda$-ECLIPSE  supports multi-concept personalization and provides additional controls (i.e., edge map) for image generation.
-In this work, we show that it is possible to perform subject-driven T2I in the latent space of a pre-trained CLIP model without depending on the diffusion UNet models; effectively reducing the heavy resource requirements.
+In this paper, we present $\lambda$-ECLIPSE, an alternative prior-training strategy that works in the latent space of a pre-trained CLIP model without relying on the diffusion UNet models.
+$\lambda$-ECLIPSE leverages the image-text interleaved pre-training for fast and effective multi-subject-driven P-T2I.
+Through extensive experiments, we establish that $\lambda$-ECLIPSE surpasses existing baselines in composition alignment while preserving concept alignment performance, even with significantly lower resource utilization.
+$\lambda$-ECLIPSE performs multi-subject driven P-T2I with just 34M parameters and is trained on a mere 74 GPU hours.
+Additionally, $\lambda$-ECLIPSE demonstrates the unique ability to perform multi-concept interpolations.
+* indicates equal contribution, $\dagger$ corresponding author: maitreya.patel@asu.edu
+Figure 1: $\lambda$-ECLIPSE  can estimate subject-specific image embeddings while maintaining the balance between concept and composition alignment in a resource-efficient way.
 1 Introduction
 The field of text-to-image (T2I) diffusion models has recently witnessed remarkable advancements, achieving greater photorealism and enhanced adherence to textual prompts.
-This has catalyzed the emergence of diverse applications, notably in controllable T2I models and personalized T2I (P-T2I) models.
-P-T2I, in particular, encompasses the intricate task of learning and reproducing novel visual concepts or subjects in varied contexts, beyond mere replication of reference images.
-This personalization task is challenging, necessitating a nuanced balance between concept and compositional alignment.
+This has catalyzed the emergence of diverse applications, notably subject-driven personalized T2I (P-T2I) models.
+In particular, this encompasses the intricate task of learning and reproducing novel visual concepts or subjects in varied contexts requiring high concept and compositional alignment.
 The complexity escalates further when multi-subject personalization is desired.
-Early works in this field, such as Textual Inversion [12] and DreamBooth [38], employed per-concept optimization strategies involving fine-tuning certain parameters within T2I diffusion models.
-However, these methods struggle with generalization challenges and are time-intensive.
-Contemporary research is pivoting towards fast personalization techniques that involve training hypernetworks and integrating new layers or parameters within pre-trained diffusion UNet models.
-As summarized in Table 1, despite their relative parameter efficiency, these methods are still resource-intensive (especially for multi-concept personalization), largely due to their reliance on the latent spaces of diffusion models.
-Additionally, controlling the inherent randomness in diffusion models remains a significant challenge, often resulting in inconsistent outputs and necessitating multiple attempts to generate a single accurate image.
-Upon further investigation, we find that most subject-driven T2I approaches build upon variants of the Latent Diffusion Model (LDM) [37], specifically Stable Diffusion models.
-LDMs employ cross-attention layers to condition diffusion models with text embeddings, necessitating a mapping of target subject images to latent spaces compatible with the diffusion models, either as text embeddings or as supplementary model parameters.
-This process involves backpropagation through the entire diffusion model, often comprising over a billion parameters, contributing to the inefficiency of existing P-T2I methods.
-In contrast, the UnCLIP [35] T2I family of LDMs offers an intriguing alternative, wherein text embeddings are mapped to vision embeddings through a diffusion transformer prior network before passing the conditioning to the diffusion UNet image generator.
-This mechanism, as illustrated in Figure 2, enables the generator to accurately reproduce original images from their visual embeddings alone.
-Consequently, our research pivots towards leveraging the UnCLIP T2I stack for P-T2I strategies.
-However, like their predecessors, these T2I diffusion priors are also parameter-heavy, with around 1 billion parameters.
-Until recently, ECLIPSE [31] showed that it is possible to compress this 1B parameter of the T2I prior model down to 33M by removing the diffusion modeling.
-Table 1: A quick overview of previous works on T2I personalization.
-Our method is the first to offer multi-concept personalization without depending on diffusion UNet models (except for inference).
-We provide the extended overview table in the supplementary materials.
-Method Multi Finetuning Diffusion Total opt. GPU
-Concepts Free Free params Hours
-Textual Inversion [12] 768 1
-DreamBooth [38] 0.9B 0.5
-ELITE [48] 77M -
-BLIP-Diffusion [21] 1.5B 2304
-IP-Adapter [50] 22M -
+Early works employed concept-specific optimization strategies involving fine-tuning certain parameters within T2I diffusion models gal2022image ; ruiz2023dreambooth ; kumari2023multi ; tumanyan2023plug ; gal2023encoder .
+Although these methods achieve state-of-the-art (SOTA) performance, they struggle with generalization and are time-intensive.
+Contemporary research is pivoting towards fast personalization techniques.
+Within this paradigm, there are two types of popular approaches: 1) Methods that involve training hypernetworks and integrating new layers or parameters within pre-trained diffusion UNet models Wei2023ELITEEV ; ye2023ip ; tewel2023key ; shah2023ziplora ; ruiz2023hyperdreambooth , and 2) MLLM-based learning of prior models that focuses on leveraging text-latent space of frozen diffusion UNet model pan2023kosmos ; sun2023generative .
+The hypernetwork-based strategy achieves single-concept customization but has not been extended to multi-concepts.
+Moreover, when combined with additional control (i.e.
+Canny edge map), they struggle to maintain the concept alignment ($\sim$30drop in performance; Section 4) and strongly favor the edge map.
+At the same time, MLLM-based approaches can perform fast multi-concept customization but require heavy computing resources.
+In Table 1, we provide the overview of various single and multi-concept customization methodologies in terms of total parameters, iterations, and GPU hours required to train the models.
+It can be observed that multi-concept customization methodologies further increase the resource requirements.
+For instance, Kosmos-G pan2023kosmos consumes 18x more resources than IP-Adapter ye2023ip .
+And Emu2 sun2023generative requires training of 19x more parameters compared to Kosmos-G.
+Hence, despite MLLMs’ seemingly useful scenarios it is not viable to blindly train them.
+Therefore, in this work, we focus on answering one question: What are the alternative methodology and design choices we can make to improve resource efficiency?
+Upon further investigation, we find that most subject-driven T2I approaches build upon variants of the Latent Diffusion Model (LDM) rombach2022high , specifically Stable Diffusion models.
+These LDMs employ cross-attention layers to condition diffusion models with text embeddings, necessitating a mapping of target subject images to latent spaces compatible with the diffusion models at the prior training stage.
+This is also known as score distillation instruction tuning for MLLMs pan2023kosmos .
+As there is no choice but to learn this text-to-image diffusion latent space, it involves backpropagation through the entire diffusion model often comprising over a billion parameters, contributing to the inefficiency of existing P-T2I methods.
+To improve the resource efficiency for multi-concept image generation, we present $\lambda$-ECLIPSE [^1] The designation $\lambda$-ECLIPSE is inspired by its conceptual alignment with the $\lambda$-calculus.
+In this context, the $\lambda$-ECLIPSE model functions similarly to a functional abstraction within $\lambda$-calculus, where it effectively binds variables.
+These variables, in our case, represent novel visual concepts that are integrated through composition prompts.
+Here, ECLIPSE indicates our architecture design choice., which leverages the properties of UnCLIP T2I models (e.g.
+DALL-E 2 ramesh2022hierarchical and Kandinsky v2.2 razzhigaev2023kandinsky ) and performs P-T2I in the compressed latent space.
+Specifically, unlike previous MLLM-based methodologies, $\lambda$-ECLIPSE aligns the output space of priors with CLIP vision space instead of the CLIP text space.
+$\lambda$-ECLIPSE takes multiple images and text instructions as input and estimates the respective vision embeddings, which can be used by the frozen diffusion UNet model from the UnCLIP stack to generate the resulting image.
+This elevates the training time dependencies on diffusion models for P-T2I; significantly contributing to the resource efficiency.
+Additionally, as diffusion or MLLM-based priors are still compute heavy due to a huge number of parameters and slower convergence, we build upon ECLIPSE patel2023eclipse and SEED ge2023making , which shows that text-to-image mapping can be optimized through contrastive pre-training.
+Here, we select ECLIPSE as preferred choice of prior architecture for best efficiency.
+At last, we propose a subject-driven instruction tuning task based on the image-text interleaved data as a pre-training strategy.
+This involves creating 2 million high-quality image-text pairs, where text embeddings linked to subjects are substituted with the respective image embeddings, which in return are considered as input to the $\lambda$-ECLIPSE.
+While $\lambda$-ECLIPSE can be plugged with these pre-trained methods, we explore the possibility of $\lambda$-ECLIPSE to incorporate Canny edge as an additional control to synergetically work with subject-driven image generation tasks.
+Figure 1 provides the overview of $\lambda$-ECLIPSE capabilities.
+Overall, we propose $\lambda$-ECLIPSE as an initial attempt to motivate future works on designing resource-efficient solutions for MLLM-based approaches.
+We summarize our main contributions as follows: 1) We introduce a training-time diffusion-independent, UnCLIP-based prior learning strategy for enabling efficient and fast multi-subject customization.
+2) Extensive experiments on Dreambench, Multibench, and ConceptBed reveal that $\lambda$-ECLIPSE (34M parameter model) trained on a mere 74 GPU hours can achieve competitive performance to that of big counterparts (having 2B-37B parameters) and improve text-composition alignment.
+3) At last, $\lambda$-ECLIPSE inherits the smooth CLIP latent space.
+This allows us to perform the seamless transition between multi-concept generated images.
+2 Related Works
+Table 1: A quick overview of previous works on P-T2I.
+Our method is the first to offer multi-concept-driven generation without depending on diffusion UNet models (except for inference).
+We provide the extended overview table in the appendix.
+Method Multi Finetuning Diffusion Total opt. Training Dataset GPU
+Concepts Free Free params Steps Size Hours
+Textual Inversion gal2022image 768 5000 - 1
+DreamBooth ruiz2023dreambooth 0.9B 800 - 0.2
+ELITE Wei2023ELITEEV 77M 135K 125K 336
+BLIP-Diffusion li2023blip 1.5B 500K 129M 2304
+IP-Adapter ye2023ip 22M 1M - 672
 […]
-Subject-Diffusion [27] 252M -
-Kosmos-G [29] 1.9B 12300
-$\lambda$-ECLIPSE (ours) 34M 74
-Figure 2: CLIP(vision) features capture the semantics and fine-grained visual details.
-UnCLIP models (e.g., Kandinsky v2.2) can effectively leverage this latent space to reconstruct the image from its embedding.
-ECLIPSE posits that text-to-image mapping can be optimized through contrastive pre-training.
-It inputs text embeddings and estimates corresponding image embeddings, ensuring strong alignment with textual features.
-Building on these insights, to enhance this framework and deepen the comprehension of novel visual subjects, we introduce a novel pre-training strategy and propose the $\lambda$-ECLIPSE  T2I prior model for the UnCLIP framework.
-Specifically, we propose a subject-driven instruction tuning task involving the image-text interleaved data.
-This involves creating 1.6 million high-quality image-text pairs, where text tokens linked to subjects are substituted with image embeddings.
-Subsequently, $\lambda$-ECLIPSE  is trained to estimate image embeddings that not only harmonize with text semantics but also encapsulate subject representations.
-This model facilitates multi-subject-driven image generation using UnCLIP models without imposing excessive resource demands.
-$\lambda$-ECLIPSE  bypasses the reliance on diffusion latent space and operates within a pre-trained CLIP latent space [33].
-We leverage Kandinsky v2.2 diffusion UNet model’s [36] capacity to interpret image structures from CLIP embeddings (refer to Figure 2).
-Importantly, CLIP image embeddings even accurately capture the details of the canny edge map [6].
-Therefore, we extend $\lambda$-ECLIPSE to incorporate canny edge maps as auxiliary guides for subject-driven T2I.
-Our extensive experimental validation demonstrates that $\lambda$-ECLIPSE  achieves unparalleled performance in single and multi-subject composition alignment while preserving the integrity of reference subjects, competitively rivaling Multimodal Large Language Models (MLLM) based methods that require intense computing resources.
-At last, we find that $\lambda$-ECLIPSE inherits the continuous latent space from the CLIP; for the first time allowing the interpolations between multiple concepts for P-T2I models.
-Figure 1 illustrates the qualitative capabilities of $\lambda$-ECLIPSE.
-In summary, our contributions are: (1) we introduce a straightforward and yet very effective diffusion-free multi-subject personalization strategy by estimating subject-aligned image embeddings in CLIP(vision) latent space, (2) despite significantly reducing resource requirements, our method achieves state-of-the-art performance, and (3) we demonstrate the potential of utilizing diffusion image generators as standalone pre-trained entities, enabling controlled outputs without necessitating explicit fine-tuning.
-2 Related Works
+Subject-Diffusion ma2023subject 252M 300K 76M -
+Kosmos-G pan2023kosmos 1.9B 800K 200M 12300
+Emu-2 sun2023generative 37B 70K 162M -
+$\lambda$-ECLIPSE (ours) 34M 100K 2M 74
+Text-to-Image Generative Models.
+Pioneering efforts in image generation, notably DALL-E ramesh2021zero and CogView ding2021cogview , leveraged autoregressive models to achieve significant results.
+Recent advancements predominantly feature diffusion models, acclaimed for their high image fidelity and diversity in text-to-image (T2I) generation.
+A notable example is Stable Diffusion, which builds upon the Latent Diffusion Model (LDM) rombach2022high and excels in semantic and conceptual understanding by transitioning training to latent space.
+Imagen saharia2022photorealistic , Pixart-$\alpha$ chen2023pixart , and DALL-E 3 betker2023improving propose using a large T5 language model to improve language understanding.
+DALL-E 2 ramesh2022hierarchical along with its UnCLIP variation models such as Kandinsky razzhigaev2023kandinsky and Karlo kakaobrain2022karlo-v1-alpha , uses a diffusion prior and diffusion UNet modules to generate images using the pre-trained CLIP radford2021learning model.
+Personalized T2I Methods.
+Approaches like Textual Inversion gal2022image , DreamBooth ruiz2023dreambooth , and Custom Diffusion kumari2023multi focus on training specific parameters to encapsulate visual concepts.
+LoRA hu2021lora and Perfusion tewel2023key target efficient fine-tuning adjustments, particularly rank 1 modifications.
+However, these methods are constrained by their requirement for concept-specific tuning.
+ELITE Wei2023ELITEEV was the first approach addressing fast customized generation for single-subject T2I.
+BLIP-Diffusion li2023blip adapts the BLIP2 encoder li2023blip2 , training approximately 1.5B parameters to enable zero-shot, subject-driven image generation.
+IP-Adapter introduces a decoupled cross-attention mechanism, negating the need to train the foundational UNet model by permitting fine-tuning of a reduced number of 22M parameters.
+Mix-of-Show gu2023mix and Zip-LoRA shah2023ziplora train individual concepts and then combine them to generate multiple subjects.
+Break-A-Scene avrahami2023break shows multi-concept capability but requires single images containing diverse objects.
+Subject Diffusion ma2023subject creates a high-quality dataset and presents the precision control for fast personalized multi-subject image generation.
+Kosmos-G and Emu2 sun2023generative , akin to Subject-Diffusion ma2023subject , employs a Multimodal Large Language Model (MLLM) for text-image embedding alignment, though it necessitates extensive parameter optimization (1.9B-37B).
+These multi-subject P-T2I methods are not only demanding in terms of parameters but also depend on a massive number of frozen parameters of the diffusion UNet model, increasing training computational loads.
+In contrast, our model, $\lambda$-ECLIPSE, forgoes test-time fine-tuning and training-time reliance on the diffusion UNet model for single and multi-concept, control-guided P-T2I, positioning it as a resource-efficient solution.
+At last, methods like GLIGEN li2023gligen , ControlNet zhang2023adding , and UniControl qin2023unicontrol incorporate additional controls (i.e., edge map, depth, segmentations) into the diffusion model to generate the desired images.
+BLIP-Diffusion, IP-Adapter, and Kosmos-G can leverage such pre-trained controls.
+However, in many scenarios, these controls are too strong, making generated images lose subject-specific details.
+We show that $\lambda$-ECLIPSE learns to balance the edge map, subjects, and composition.
+We offer a more comprehensive review of related works in the appendix.
+3 Method
 Figure 2: This figure illustrates the three stages of the $\lambda$-ECLIPSE pipeline.
-1) Create the image-text interleaved features by utilizing the pre-trained CLIP encoders.
+1) Create the image-text interleaved features using frozen CLIP.
 2) Pre-train the $\lambda$-ECLIPSE (34M parameters) using Eq.
 1, which ensures the mapping to the desired latent space given the image-text interleaved data.
-During inference, we just need a single-forward pass to estimate the corresponding image embedding.
-3) During inference, the Kandinsky v2.2 diffusion UNet model takes the output from the $\lambda$-ECLIPSE (prior) and generates the corresponding images.
-Text-to-Image Generative Models.
-Pioneering efforts in image generation, notably DALL-E [34] and CogView [11], leveraged autoregressive models to achieve significant results.
-Recent advancements predominantly feature diffusion models, acclaimed for their high image fidelity and diversity in text-to-image (T2I) generation.
-A notable example is Stable Diffusion, which builds upon the Latent Diffusion Model (LDM) [37] and excels in semantic and conceptual understanding by transitioning training to latent space.
-Imagen [40], Pixart-$\alpha$ [8], and DALL-E 3 [5] propose using a large T5 language model to improve language understanding.
-DALL-E 2 [35] along with its UnCLIP variation models such as Kandinsky [36] and Karlo, uses a diffusion prior and diffusion UNet modules to generate images using the pre-trained CLIP [33] model.
-Personalized T2I Methods.
-Approaches like Textual Inversion [12], DreamBooth [38], and Custom Diffusion [20] focus on training specific parameters to encapsulate visual concepts.
-LoRA [16] and Perfusion [45] target efficient fine-tuning adjustments, particularly rank 1 modifications.
-However, these methods are constrained by their requirement for concept-specific tuning.
-ELITE [48] was the first approach addressing fast personalization for single-subject T2I.
-BLIP-Diffusion [21] adapts the BLIP2 encoder [22], training approximately 1.5B parameters to enable zero-shot, subject-driven image generation.
-IP-Adapter [50] introduces a decoupled cross-attention mechanism, negating the need to train the foundational UNet model by permitting fine-tuning of a reduced parameter set (22M).
-Mix-of-Show [14] and Zip-LoRA [42] train individual concepts and then combine them to generate multiple subjects.
-Break-A-Scene [4] shows multi-concept capability but requires single images containing diverse objects.
-Subject Diffusion [27] creates a high-quality dataset and presents the precision control for fast multi-subject image generation.
-Kosmos-G [29], akin to Subject-Diffusion, employs a Multimodal Large Language Model (MLLM) for text-image embedding alignment, though it necessitates extensive parameter optimization (1.9B).
-These multi-subject P-T2I methods are not only demanding in terms of parameters but also depend on a massive number of frozen parameters of the diffusion UNet model, increasing training computational loads.
-In contrast, our model, $\lambda$-ECLIPSE, forgoes test-time fine-tuning and training-time reliance on the diffusion UNet model for single and multi-concept, control-guided P-T2I, positioning it as a resource-efficient solution.
-At last, methods like GLIGEN [23], ControlNet [51], and UniControl [32] incorporate additional controls (i.e., edge map, depth, segmentations) into the diffusion model to generate the desired images.
-BLIP-Diffusion, IP-Adapter, and Kosmos-G can leverage such pre-trained controls.
-However, in many scenarios, these controls are too strong which makes generated images lose subject-specific details.
-While $\lambda$-ECLIPSE can be plugged with these pre-trained methods, we explore the possibility of learning canny edge as the additional guidance for P-T2I.
-We show that $\lambda$-ECLIPSE learns to balance the edge map, subjects, and text composition.
-3 Method
-In this section, we introduce $\lambda$-ECLIPSE, our approach to multi-subject text-to-image personalization.
-This method extends the capabilities of the existing ECLIPSE  framework through a novel image-text interleaved pretraining strategy, notably omitting the need for explicit diffusion modeling.
+3) During inference, the frozen Kandinsky v2.2 diffusion UNet model takes the output from the $\lambda$-ECLIPSE  and generates the image.
+In this section, we introduce $\lambda$-ECLIPSE, our approach to multi-subject personalized text-to-image generation.
+Our method combines the contrastive text-to-image strategy from ECLIPSE with the novel image-text interleaved pretraining strategy, notably omitting the need for explicit diffusion modeling.
 Our approach mainly capitalizes on the efficient utilization of the CLIP latent space.
 Figure 2 outlines the end-to-end framework.
-The primary objective of $\lambda$-ECLIPSE is to facilitate both single and multi-subject T2I generation processes, accommodating controlled conditional guidance such as edge maps.
-Initially, we detail the problem formulation and elaborate on the extended design of the ECLIPSE  T2I prior module.
+The primary objective of $\lambda$-ECLIPSE is to facilitate single and multi-subject P-T2I generation processes, accommodating edge maps as conditional guidance.
+Initially, we detail the problem formulation and elaborate on the UnCLIP stack design of the $\lambda$-ECLIPSE.
 Subsequently, we delve into the image-text interleaved training methodology.
-This fine-tuning process enables the ECLIPSE  model to harness semantic correlations between CLIP image and text latent spaces while preserving subject-specific visual features.
-Upon attaining the target embedding, the diffusion UNet model, derived from Kandinsky v2.2, is employed to reconstruct the image.
+This fine-tuning process enables the $\lambda$-ECLIPSE to harness semantic correlations between CLIP image and text latent spaces while preserving subject-specific visual features.
 3.1 Text-to-Image Prior Mapping
+Figure 3: CLIP(vision) features capture the semantics and fine-grained visual details.
+Each input is given as input to the Kandinsky v2.2 and re-generated from the decoder.
+(Top: Real-images, Bottom: Canny edge)
 In the UnCLIP T2I models, the objective of the text-to-image prior model ($f_{\theta}$) is to establish a proficient text-to-image embedding mapping.
@@
 The visual embedding predictions ($\hat{z_{x}}=f_{\theta}(z_{y})$) are then effectively utilized by the diffusion image generators ($h_{\phi}$), which are inherently conditioned on these vision embeddings ($h_{\phi}:z_{x}\rightarrow x$).
-As illustrated in Figure 2, these diffusion image generators possess a remarkable capability for precise image reproduction, since $z_{x}$ encompasses comprehensive image information.
+In our experiments, we utilize the Kandinsky v2.2 diffusion UNet model as $h_{\phi}$.
+As shown in Figure 3, the CLIP vision encoder preserves the necessary features in $z_{x}$ required to reconstruct the input image and achieves a similar concept alignment score (DINO: 0.66) as finetuning-based DreamBooth method ruiz2023dreambooth .
 Our goal is to accurately estimate the image embedding $\hat{z}_{x}$, incorporating the subject representations, thereby eliminating reliance on $h_{\phi}$ during training.
 Existing LDM-based P-T2I methods are limited by the LDM’s singular module approach ($h_{\phi}:z_{y}\rightarrow x$).
-Consequently, mastering the latent space of $h_{\phi}$ becomes essential for effective P-T2I for the baseline methodologies.
-We propose a mapping function, $f^{\prime}_{\theta}$, which processes text representations ($z_{y}$) alongside subject ($x_{i}$) specific representations ($z_{x_{i}}$), to derive an image embedding that encapsulates both text prompts and subject visuals ($\hat{z}_{x}$).
-The challenge lies in harmonizing $z_{x_{i}}$ and $z_{y}$ within $f^{\prime}_{\theta}:(z_{y},z_{x_{i}})\rightarrow\hat{z}_{x}$, ensuring alignment while preventing overemphasis on either aspect, as this could compromise composition alignment.
-To address this, we employ the contrastive pre-training strategy after [31]:
-$\begin{split}\mathcal{L}_{prior}=&\underset{\begin{subarray}{c}\epsilon\sim\mathcal{N}(0,I)\\ z_{y},z_{x}\end{subarray}}{\operatorname{\mathbb{E}}}\Big{[}||z_{x}-f_{\theta}(\epsilon,z_{y})||_{2}^{2}\Big{]}\\ &-\frac{\lambda}{N}\sum_{i=0}^{N}\log\frac{\exp(\langle\hat{z}^{i}_{x},z^{i}_{y}\rangle/\tau)}{\sum_{j\in[N]}\exp(\langle\hat{z}^{i}_{x},z^{j}_{y}\rangle/\tau)}.\end{split}$ (1)
-Here, $\lambda$ serves as the hyperparameter.
-The first loss term measures the mean-squared error between the estimated and actual image embeddings, primarily ensuring concept alignment.
-However, our preliminary studies reveal that exclusive reliance on this term diminishes composition alignment, echoing observations made by [31].
-Therefore, we stick with the contrastive loss component to bolster compositional generalization, with $\lambda$ balancing concept and composition alignment.
-Figure 4: Qualitative examples of $\lambda$-ECLIPSE.
-Figure 5: This figure illustrates the qualitative comparison of $\lambda$-ECLIPSE w.r.t.
-the state-of-the-art approaches for single-subject T2I generations.
-We choose concepts and prompts from the Dreambench dataset.
-For all methods, concepts, and prompts, we generate the four images and select the best one that represents the concept and composition more accurately.
+Consequently, mastering the latent space of $h_{\phi}$ becomes essential for effective P-T2I for the baseline methodologies, which limits the previous methodologies.
+We propose a new mapping function, $f_{\theta}$, which processes text representations ($z_{y}$) alongside subject ($x_{k}$) specific visual representations ($z_{x_{k}}$), to derive an image embedding that encapsulates both text prompts and subject visuals ($\hat{z}_{x}$).
+The challenge lies in harmonizing $z_{x_{k}}$ and $z_{y}$ within $f_{\theta}:(z_{y},z_{x_{k}})\rightarrow\hat{z}_{x}$, ensuring alignment while preventing overemphasis on either aspect, as this could compromise composition alignment.
+To address this, we employ the contrastive pre-training strategy after patel2023eclipse :
+$\begin{split}\mathcal{L}_{prior}=\underset{\begin{subarray}{c}\epsilon\sim\mathcal{N}(0,I)\\ z_{y},z_{x_{k}}\end{subarray}}{\mathbb{E}}\Big{[}||z_{x}-f_{\theta}(\epsilon,z_{y},z_{x_{k}})||_{2}^{2}\Big{]}-\frac{\lambda}{N}\sum_{i=0}^{N}\log\frac{\exp(\langle\hat{z}^{i}_{x},z^{i}_{y}\rangle/\tau)}{\sum_{j\in[N]}\exp(\langle\hat{z}^{i}_{x},z^{j}_{y}\rangle/\tau)}.\end{split}$ (1)
+[…]
 Additional Control-based T2I Prior Mapping.
@@
 Consequently, we have augmented $\lambda$-ECLIPSE to accommodate an additional modality, a Canny edge map, providing more refined control over subject-driven image generation.
-This entails modifying the prior model to accept additional conditions ($f_{\theta}^{\prime}:(z_{y},z_{x_{i}},z_{c})\rightarrow\hat{z}_{x}$, where $z_{c}$ symbolizes the additional modality embedding).
+This entails modifying the prior model to accept additional conditions ($f_{\theta}^{\prime}:(z_{y},z_{x_{i}},z_{c})\rightarrow\hat{z}_{x}$, here $z_{c}$ symbolizes the additional modality embedding).
 During training, we drop $z_{c}$ for 1% to improve the unconditional generations.
-Incorporating canny edge maps during training enhances stability and broadens the generalization capabilities of $\lambda$-ECLIPSE, yielding benefits even in the absence of these controls during inference.
-Our findings indicate that $\lambda$-ECLIPSE successfully learns a unified mapping function, effectively estimating target image representations from a synergy of text, subjects, and edge map.
+This enhances the stability and broadens the generalization capabilities of $\lambda$-ECLIPSE, yielding benefits even in the absence of these controls during inference.
+Our results demonstrate that $\lambda$-ECLIPSE  learns a unified mapping function, accurately estimating target image representations through the effective integration of text, image, and edge maps.
 3.2 Image-text Interleaved Training
-Our approach targets developing a versatile prior model capable of processing diverse inputs to estimate visual outputs.
+Figure 4: This figure illustrates a qualitative comparison of $\lambda$-ECLIPSE  with contemporary approaches for single-subject T2I generations, utilizing concepts and prompts from the Dreambench dataset.
+For each method, concept, and prompt, we generate four images and select the one that most accurately represents the queried concept and composition.
+Our approach targets developing a versatile prior model capable of processing diverse inputs to estimate target visual outputs.
 Drawing from earlier methodologies, a straightforward solution involves concatenating different inputs, like combining text (“a dog wearing sunglasses”) with respective concept-specific images.
-However, preliminary experiments indicated that this method does not effectively capture the intricate relationships between target text tokens (e.g., “dog”) and the corresponding concept images.
+Preliminary experiments indicated that this method does not effectively capture the intricate relationships between target text tokens (e.g.
+“dog”) and the corresponding concept images, especially when multiple concepts are present.
 To address this, we adopt the interleaved pre-training strategy used in Kosmos-G, but with a notable modification to enhance resource efficiency.
-We incorporate pretrained CLIP text and vision encoders for extracting modality-specific embeddings—separating text-only from subject-specific images.
-The key refinement in our process is the substitution of subject token embeddings with corresponding vision embeddings instead of introducing additional tokens to handle the image embeddings via resampler [2].
-This alteration allows us to bypass the need to train the prior or MLLMs, significantly improving the model’s proficiency in handling interleaved data.
-For the generation of high-quality training datasets, we carefully selected 2 million high-quality images from the LAION dataset [41], each with a resolution of 1024x1024.
-Utilizing BLIP2, we generate captions for these images and employ SAM [19] for extracting noun or subject-specific segmentation masks.
+We incorporate pretrained frozen CLIP text and vision encoders for extracting modality-specific embeddings—separating text-only from subject-specific images.
+The key refinement in our process is the substitution of subject token-specific text embeddings with corresponding vision embeddings instead of introducing additional trainable tokens to handle the image embeddings via resampler alayrac2022flamingo .
+First, we extract reference concept visual features ($z_{x_{k}}\in\mathcal{R}^{1\mathrm{x}1280}$) from the CLIP vision encoder.
+Similarly, we also extract the text prompt features ($z_{y}\in\mathcal{R}^{77\mathrm{x}1280}$) from the last layer of the CLIP text encoder.
+Here, 1280 is the CLIP-specific feature dimension.
+At last, we replace the concept noun corresponding latent features from $z_{y}$ with $z_{x_{k}}$; resulting in image-text interleaved features while preserving the contextual information of the text features.
+This alteration allows us to bypass the need to train the big priors models (e.g.
+MLLMs), significantly improving the model’s proficiency in handling interleaved data.
+For the generation of high-quality training datasets, we carefully selected 2 million high-quality images from the LAION dataset schuhmann2022laion , each with a resolution of 1024x1024.
+Utilizing BLIP2, we generate captions for these images and employ SAM kirillov2023segment for extracting noun or subject-specific segmentation masks.
 Given the CLIP model’s requirement for 224x224 resolution images, we avoid resizing the masks within their original resolutions.
-Instead, we opt for cropping the area of interest using DINO [24], followed by resizing the masked object while preserving its aspect ratio.
+Instead, we opt for cropping the area of interest using Grounding DINO liu2023grounding , followed by resizing the masked object while preserving its aspect ratio.
 This technique is crucial in retaining maximum visual information for each subject during the training phase.
 We provide more details about the filters used in the supplementary material.
-In summary, the ECLIPSE  prior, trained with our image-text interleaved data and supplementary conditions, presents an efficient pathway for P-T2I.
-The culmination of this process yields the $\lambda$-ECLIPSE model, a refined version characterized by enhanced effectiveness in handling and interpreting multimodal data.
+3.3 Additional Finetuning
+Due to the nature of UnCLIP models, even if $\lambda$-ECLIPSE is very accurate, the diffusion UNet model ($h_{\phi}$) may not be effective in generating very unique visual representations.
+However, such behavior is common across the fast P-T2I methods and they lack in terms of maintaining performance compared to the finetuning-based methods (as outlined in Table 2).
+Therefore, we extend the $\lambda$-ECLIPSE and perform concept-specific finetuning.
+Compared to the traditional finetuning methodologies (such as DreamBooth), $\lambda$-ECLIPSE provides very unique advantages.
+As $\lambda$-ECLIPSE prior model ($f_{\theta}$) is pre-trained for personalization, there is no need for further finetuning the $f_{\theta}$ and we need to only finetune diffusion UNet model.
+Importantly, the fine-tuning of the $h_{\phi}$ does not depend on the text embeddings ($z_{y}$).
+Hence, this leads to stable fine-tuning of the $h_{\phi}$; unlike DreamBooth on stable diffusion that observes catastrophic forgetting (Section D).
+The new fine-tuning objective is:
+$\mathcal{L}_{decoder}=\underset{\begin{subarray}{c}\epsilon\sim N(0,I)\\ t\sim[0,T],(z_{x})\end{subarray}}{\mathbb{E}}\Big{[}||\epsilon-h_{\phi}(x^{(t)},t,z_{x})||_{2}^{2}\Big{]}.$ (2)
+Here, $z_{x}$ is the visual feature of the reference concept image $x$.
+Notably, we do not need to use regularization from the DreamBooth as text alignment is already ensured during the pretraining stage of $\lambda$-ECLIPSE.
+Moreover, this finetuning can be performed across the set of given visual concepts altogether in a single model without degrading performance.
+In summary, the prior model, trained with our image-text interleaved data and supplementary condition, presents an efficient pathway for resource-efficient multi-subject-driven image generations.
 4 Experiments
-In this section, we first introduce the experimental setup, evaluations, and baselines. Later, we delve into the qualitative and quantitative results.
+In this section, we first introduce the experimental and evaluation setups. Later, we delve into the qualitative and quantitative results.
 Training and inference details.
@@
 We drop the Canny edge map during inference unless specified explicitly.
+Unless specified all results (quantitative and qualitative) are without concept-specific additional fine-tuning.
 Evaluation setup.
-We primarily utilize Dreambench (encompassing 30 unique concepts with 25 prompts per concept) for qualitative and quantitative evaluations using DINO and CLIP-based metrics [38].
-Due to their limitations, we extend our evaluations on the ConceptBed [30] benchmark (covering 80 diverse imagenet concepts and a total of 33K composite prompts), where we report performance on concept replication, concept, and composition alignment using the Concept Confidence Deviation ($CCD(\downarrow)$) metric.
-4.1 Results & Analysis
+We primarily utilize Dreambench (encompassing 30 unique concepts with 25 prompts per concept) for qualitative and quantitative evaluations using DINO and CLIP-based metrics ruiz2023dreambooth .
+Due to their limitations, we extend our evaluations on the ConceptBed patel2023conceptbed benchmark (covering 80 diverse imagenet concepts and a total of 33K composite prompts), where we report performance on concept replication, concept, and composition alignment using the Concept Confidence Deviation ($CCD$) metric patel2023conceptbed .
+We extend Dreambench for multi-subject customization and present the Multibench dataset.
+Multibench contains about 24 unique concepts and 15 diverse prompts that result in 904 two-subject specific prompts and 1476 three-subject specific prompts.
+We provide further details about the Multibench in supplementary materials.
 Table 2: Quantitative comparisons of different methodologies on Dreambench.
 The Bold and underline represent the metric-specific first and second-ranked methods, respectively.
-* represents that we reproduce the open-source methodologies.
-Other results are borrowed from the Subject-Diffusion [27].
-Method Base Model DINO (normal-$\uparrow$) CLIP-I (normal-$\uparrow$) CLIP-T (normal-$\uparrow$)
-Textual Inversion SDv1.5 0.569 0.780 0.255
-DreamBooth SDv1.5 0.668 0.803 0.305
-Custom Diffusion SDv1.5 0.643 0.790 0.305
-Re Imagen Imagen 0.600 0.740 0.270
-ELITE SDv1.4 0.621 0.771 0.293
-Subject-Diffusion SDv1.5 0.711 0.787 0.293
-BLIP-Diffusion* SDv1.5 0.603 0.793 0.291
-IP-Adapter* SDv1.5 0.629 0.827 0.264
-IP-Adapter* SDXL 0.613 0.810 0.292
-Kosmos-G* SDv1.5 0.618 0.822 0.250
+* represents that we re-benchmark the performance from open-source weights.
+Method Base Model Params GPU Hours DINO (normal-$\uparrow$) CLIP-I (normal-$\uparrow$) CLIP-T (normal-$\uparrow$)
+Textual Inversion SDv1.5 768 1 0.569 0.780 0.255
+DreamBooth SDv1.5 0.9B 0.2 0.668 0.803 0.305
+Custom Diffusion SDv1.5 57M 0.2 0.643 0.790 0.305
+BLIP-Diffusion SDv1.5 0.9B 0.1 0.670 0.805 0.302
 […]
-Figure 6: Multi-subject qualitative examples illustrating the comparison between $\lambda$-ECLIPSE and other SOTA baselines.
+Re-Imagen Imagen - - 0.600 0.740 0.270
+ELITE SDv1.4 77M 336 0.621 0.771 0.293
+Subject-Diffusion SDv1.5 252M - 0.711 0.787 0.293
+BLIP-Diffusion* SDv1.5 1.5B 2304 0.603 0.793 0.291
+IP-Adapter* SDv1.5 22M 672 0.629 0.827 0.264
+IP-Adapter* SDXL 22M 672 0.613 0.810 0.292
+[…]
+Emu2* SDXL 37B - 0.563 0.765 0.273
+$\lambda$-ECLIPSE* Kv2.2 34M 74 0.613 0.783 0.307
 Table 3: Quantitative comparisons of different methodologies on ConceptBed.
@@
 The Bold and underline represent the metric-specific first and second-ranked methods, respectively.
-* represents that we reproduce the open-source methodologies.
-Other results are borrowed from the ConceptBed [30].
+* represents our benchmarking.
 Method Base Concept Concept Composition
-Model Replication Alignment Alignment
+Model Replication (normal-$\downarrow$) Alignment (normal-$\downarrow$) Alignment (normal-$\downarrow$)
 Textual Inversion SDv1.4 0.0662 0.1163 0.1436
@@
 […]
-Figure 7: Qualitative examples for edge-guided P-T2I showcasing the comparison between BLIP-Diffusion vs. $\lambda$-ECLIPSE.
+4.1 Results & Analysis
+Figure 5: Qualitative results categorized by generative capabilities.
 Quantitative comparison.
-The quantitative assessments are detailed in Table 2 and Table 3, focusing on the single-concept T2I task.
-For Dreambench, we generate and evaluate four images per prompt, reporting average performance on three metrics (DINO, CLIP-I, and CLIP-T).
-In the case of ConceptBed, we process each of the 33K prompts to generate a single image.
-The results, as depicted in these tables, highlight $\lambda$-ECLIPSE’s superior performance in composition tasks and its competitive edge over even fine-tuning-based methodologies, while maintaining robust concept alignment.
-$\lambda$-ECLIPSE exhibits a notable proficiency in concept replication, albeit with a marginal trade-off in concept alignment for enhanced composition fidelity.
-Comparatively, IP-Adapter, ELITE, and BLIP-Diffusion prioritize concept alignment, often at the expense of composition alignment.
-Noteworthy is $\lambda$-ECLIPSE’s efficiency, achieved with significantly fewer resources.
+The quantitative assessments detailed in Table 2 and Table 3 focus on the single-concept T2I task, while Table 4 shows the results on multi-concept-driven image generation.
+For Dreambench and Multibench, we generate and evaluate four images per prompt, reporting average performance on three metrics (DINO, CLIP-I, and CLIP-T).
+In the case of ConceptBed, we process each of the 33K prompts to generate a single concept image.
+The results, as depicted in these tables, highlight $\lambda$-ECLIPSE’s superior performance in composition alignment while maintaining competitive concept alignment.
+Analysis on ConceptBed (Table 3) indicates that $\lambda$-ECLIPSE exhibits a notable proficiency in concept replication, albeit with a marginal trade-off in concept alignment for enhanced composition fidelity.
+Comparatively, all baselines prioritize concept alignment, often at the expense of composition alignment.
+While $\lambda$-ECLIPSE improves the CLIP-T while preserving the DINO; achieved with significantly fewer resources.
+However, a significant gap remains between finetuning-based and fast P-T2I methods.
+We further perform concept-specific fine-tuning (as described in Section 3.3).
+As shown in Table 2, $\lambda$-ECLIPSE outperforms the DreamBooth and BLIP-Diffusion in terms of concept alignment (DINO) while maintaining the performance on composition alignment (CLIP-T).
+Notably, Multibench results (Table 4) indicate that $\lambda$-ECLIPSE significantly outperforms the Kosmos-G (2B params) and Emu2 (37B params) in terms of CLIP-T while maintaining the DINO performance.
+Therefore, we can conclude that $\lambda$-ECLIPSE is the most resource-efficient compared, especially when compared to the MLLM-based methods.
+Table 4: Quantitative comparisons of different methodologies on Multibench. The Bold and underline represent the metric-specific first and second-ranked methods on each metric, respectively.
+Two Subjects Three Subjects
+Kosmos-G Emu2 $\lambda$-ECLIPSE Emu2 $\lambda$-ECLIPSE
+DINO (normal-$\uparrow$) 0.4549 0.4451 0.4478 0.3168 0.3420
+CLIP-I (normal-$\uparrow$) 0.7759 0.7397 0.7409 0.6231 0.6463
+CLIP-T (normal-$\uparrow$) 0.2493 0.2673 0.3327 0.2819 0.3469
+Table 5: Quantitative results of Canny edge controlled P-T2I of different methodologies on Dreambench.
+The Bold and underline represent the metric-specific first and second-ranked methods, respectively.
+Red highlighted numbers indicate the relative percentage drop for concept alignment compared to Table 2.
+Method DINO (normal-$\uparrow$) CLIP-I (normal-$\uparrow$) CLIP-T (normal-$\uparrow$)
+BLIP-Diffusion* 0.423429.7% 0.7119 0.3152
+IP-Adapter* 0.428131.9% 0.7315 0.3034
+[…]
 Qualitative comparisons.
-In Figure 5, we present a range of single subject-specific images generated by various methodologies including BLIP-Diffusion, IP-Adapter, Kosmos-G, and $\lambda$-ECLIPSE.
+In Figure 4, we present a range of single subject-specific images generated by various methodologies including BLIP-Diffusion, IP-Adapter, Kosmos-G, Emu2, and $\lambda$-ECLIPSE.
 $\lambda$-ECLIPSE demonstrates exemplary proficiency in composition while ensuring concept alignment.
-In contrast, the baselines often overemphasize reference images (notably in rows 2 and 4) or exhibit concept dilution (as seen in rows 6 and 8), leading to higher concept alignment but compromised composition.
-Additionally, Figure 6 exhibits $\lambda$-ECLIPSE’s multi-concept generation prowess, in comparison to ZipLoRA (fine-tuning-based approach) and Kosmos-G (Multimodal LLM-based approach), underscoring its capability to rival compute-intensive methodologies.
-Figure 2 previously demonstrated Kandinsky v2.2’s ability to partially reconstruct canny edge maps while preserving semantic integrity.
-Building on this, Figure 7 offers qualitative insights, comparing BLIP-Diffusion and $\lambda$-ECLIPSE in edge-guided subject-driven T2I scenarios.
-Here, BLIP-Diffusion adheres strictly to the imposed conditions, often at the cost of concept retention (rows 1, 3, and 4).
-Conversely, $\lambda$-ECLIPSE not only complies with these conditions but also adeptly balances the need for high concept and composition alignment.
-Figure 8: Interpolation between four concepts.
-Here, we estimate the image embedding using $\lambda$-ECLIPSE corresponding to each corner and then interpolate from top to bottom and left to right.
-At last, we use the Kandinsky v2.2 diffusion UNet model to generate the images with fixed random seeds from these sets of image embeddings.
+In contrast, the baselines often overemphasize reference images or exhibit concept dilution, leading to higher concept alignment but compromised composition.
+Interestingly, we find that Emu2 can capture the single-subjects but it fails to reproduce them with complex text compositions (as shown in Figure 4).
+Similarly, Figure 5(a) exhibits $\lambda$-ECLIPSE’s multi-concept generation prowess, in comparison to ZipLoRA (fine-tuning-based approach) along with Kosmos-G and Emu2 (Multimodal LLM-based approaches), underscoring its capability to rival compute-intensive methods.
+We discuss additional examples and limitations in the appendix.
+That said, even though $\lambda$-ECLIPSE improves the performance over the baselines, this is still not enough and it signifies the challenges associated with fast multi-concept personalization.
+(a) Multi-subject qualitative examples.
+(b) Qualitative examples for edge-guided P-T2I.
+Figure 6: Qualitative comparison between $\lambda$-ECLIPSE and other baselines.
+Canny edge controlled image generation.
+As shown in Figure 5(b), the baseline (BLIP-Diffusion) adheres strictly to the imposed edge maps, often at the cost of concept retention (rows 1, 3, and 4).
+This leads to a large number of unwanted artifacts in the generated images.
+To further ground this behavior, we first generated images using Stable Diffusion v1.5 for Dreambench prompts without customization then we extracted the Canny edge map and used this edge map to control the subject-driven image generations.
+At last, we report the performance in Table 5.
+It can be observed that both baselines IP-Adapter and BLIP-Diffusion drop the DINO score by 30%, which follows the qualitative results.
+While $\lambda$-ECLIPSE do not follow the Canny edge precisely but preserves the concept alignment and improves the performance relatively by 21%.
 Table 6: Ablation studies w.r.t.
@@
 Model Concept Composition
-Alignment Alignment
-Projection loss 0.394 0.008
-w/ contrastive loss (lambda=0.5) 0.435 -0.043
-w/ contrastive loss (lambda=0.2) 0.402 -0.026
+Alignment (normal-$\downarrow$) Alignment (normal-$\downarrow$)
+Projection loss (i.e. $\lambda$=0.0) 0.394 0.008
+ w/ contrastive loss ($\lambda$=0.5) 0.435 -0.043
+ w/ contrastive loss ($\lambda$=0.2) 0.402 -0.026
 […]
+Figure 7: Interpolation between four concepts.
+Here, we estimate the image embedding using $\lambda$-ECLIPSE corresponding to each corner and then interpolate from top to bottom and left to right.
+At last, we use the Kandinsky v2.2 diffusion UNet model to generate the images with fixed random seeds from these sets of image embeddings.
 Ablations.
@@
 A comparative analysis of these baselines is conducted against the fully equipped $\lambda$-ECLIPSE model, which incorporates $\mathcal{L}_{prior}$ (Eq.1) with $\lambda=0.2$ and utilizes Canny edge maps during training.
-As Table 4 illustrates, relying solely on projection loss results in high concept alignment but compromises compositions.
+Relying solely on projection loss results in high concept alignment but compromises compositions (Table 6).
 The contrastive loss variant with $\lambda=0.5$ enhances composition alignment at the expense of concept alignment, whereas $\lambda=0.2$ achieves a more balanced performance.
 Crucially, the integration of Canny edge maps during training optimally balances both alignments and, specifically, improves the concept alignment.
-Here, the negative values indicate that the $CCD$ oracle model performs similarly to real images having the same set of compositions.
+The negative values indicate that the $CCD$ oracle model is highly confident in the generated images.
 Multi-subject interpolation.
@@
 We selected two distinct dog breeds ($<$dog1$>$, $<$dog2$>$) and two types of hats ($<$hat1$>$, $<$hat2$>$) as the concepts.
-$\lambda$-ECLIPSE was then used to estimate image embeddings for all four possible combinations, each corresponding to the input phrase “a $<$dog$>$ wearing a $<$hat$>$.” Figure 8 showcases a gradual and seamless transition in the synthesized images from the top left to the bottom right.
-Unlike current diffusion models, which often exhibit sensitivity to input variations necessitating numerous iterations for desired outcomes, $\lambda$-ECLIPSE inherits CLIP’s smooth latent space.
-This inheritance not only facilitates progressive changes in the conceptual domain but also extends the model’s utility by enabling personalized multi-subject interpolations, a feature that is not inherently available in previous P-T2I methods.
+$\lambda$-ECLIPSE was then used to estimate image embeddings for all four possible combinations, each corresponding to the input phrase “a $<$dog$>$ wearing a $<$hat$>$.” Fig.
+7 showcases a gradual and seamless transition in the synthesized images from the top left to the bottom right.
+Unlike current diffusion models, which often exhibit sensitivity to input variations requiring numerous iterations of user interactions for desired outcomes, $\lambda$-ECLIPSE inherits CLIP’s smooth latent space.
+This not only facilitates progressive changes in the conceptual domain but also extends the model’s utility by enabling personalized multi-subject interpolations.
 5 Conclusion
-In this paper, we have introduced a novel diffusion-free methodology for personalized text-to-image (T2I) applications, utilizing the latent space of the pre-trained CLIP model with high efficiency.
-Our $\lambda$-ECLIPSE model, trained on an image-text interleaved dataset of superior quality, marks a significant advancement in the field.
-This model outperforms existing methods by its capability to execute single-concept, multi-concept, and edge-guided controlled P-T2I tasks using a singular model framework, while simultaneously minimizing resource utilization.
-Notably, $\lambda$-ECLIPSE sets a new benchmark in achieving state-of-the-art results in composition alignment, coupled with exceptional maintenance of subject fidelity.
+In this paper, we have introduced a novel diffusion-free methodology for personalized text-to-image (P-T2I) applications, utilizing the latent space of the pre-trained CLIP model with high efficiency.
+Our $\lambda$-ECLIPSE model, trained on an image-text interleaved dataset, achieves the capability to execute single-concept, multi-concept, and edge-guided controlled P-T2I tasks using a singular model framework, while simultaneously minimizing resource utilization.
+Notably, $\lambda$-ECLIPSE sets a new benchmark in achieving competitive performance in terms of concept and composition alignment.
 Furthermore, our research illuminates the potential of $\lambda$-ECLIPSE in exploring and leveraging the smooth latent space.
 This capability unlocks new avenues for interpolating between multiple concepts and their amalgamation, thereby generating entirely novel concepts.
-Our findings underscore a promising pathway where pre-trained diffusion image generators can be effectively controlled without necessitating extra supervision.
+Our findings underscore a promising pathway to improve MLLMs to effectively control the pre-trained diffusion image models without necessitating extra supervision.
 Acknowledgments
 This work was supported by NSF RI grants #1750082, #2132724, and CPS grant #2038666.
+We thank the Research Computing (RC) at Arizona State University (ASU) for providing computing resources.
 The views and opinions of the authors expressed herein do not necessarily state or reflect those of the funding agencies and employers.
 Appendix A Preliminaries for T2I Diffusion Models
-As evidenced in numerous contemporary studies regarding T2I models, Stable Diffusion (SD) has emerged as a predominant backbone for text-to-image (T2I) models.
+As evidenced in numerous contemporary studies regarding T2I models, Stable Diffusion (SD) [38] has emerged as a predominant backbone for T2I models.
 SD involves training diffusion models in latent space, reversing a forward diffusion process that introduces noise into the image.
 A notable feature of SD is its integration of cross-attention, facilitating various conditions like text input.
-Operating in VQVAE latent space, SD not only achieves exceptional generative performance surpassing that in pixel space but also significantly reduces computational demands.
+Operating in VQ-VAE latent space, SD not only achieves exceptional generative performance surpassing that in pixel space but also significantly reduces the computational demands.
 UnCLIP models (such as DALL-E 2) are very similar to the Stable Diffusion.
-In contrast, the UnCLIP stack takes the modular approach.
+In contrast, the UnCLIP takes the modular approach.
 UnCLIP first trains the diffusion text-to-image to the image prior ($f_{\theta}$) to estimate the image embeddings ($z_{x}$) from the text embeddings ($z_{y}$).
@@
 Given the forward function $z_{x}^{(t)}\sim q(t,z_{x})$, the goal of $f_{\theta}$ is to directly estimate $z_{x}$ for all timesteps $t\sim[0,T]$ as:
-$\mathcal{L}_{prior}=\underset{\begin{subarray}{c}t\sim[0,T],\\ z_{x}^{(t)}\sim q(t,z_{x})\end{subarray}}{\operatorname{\mathbb{E}}}\Big{[}||z_{x}-f_{\theta}(z_{x}^{(t)},t,z_{y})||_{2}^{2}\Big{]}.$ (2)
-ECLIPSE presents the contrastive learning strategy (Eq.
-1) instead of minimizing Eq.
-2.
+$\mathcal{L}_{prior}=\underset{\begin{subarray}{c}t\sim[0,T],\\ z_{x}^{(t)}\sim q(t,z_{x})\end{subarray}}{\mathbb{E}}\Big{[}||z_{x}-f_{\theta}(z_{x}^{(t)},t,z_{y})||_{2}^{2}\Big{]}.$ (3)
+ECLIPSE proposes the contrastive learning strategy (Eq.
+1 – main paper) instead of minimizing Eq.
+3.
 The diffusion image generator is trained by following standard $\epsilon$-prediction formulation.
 Here, $h_{\phi}$ will estimate the ground truth added Gaussian noise $\epsilon\sim N(0,I)$, given the noise image $X^{(t)}$ for all timesteps $t\sim[0,T]$ and input conditions (such as $z_{x}$, $z_{y}$).
-$\mathcal{L}_{decoder}=\underset{\begin{subarray}{c}\epsilon\sim N(0,I)\\ t\sim[0,T],\\ (z_{x},~{}z_{y})\end{subarray}}{\operatorname{\mathbb{E}}}\Big{[}||\epsilon-h_{\phi}(x^{(t)},t,z_{x},z_{y})||_{2}^{2}\Big{]}.$ (3)
+$\mathcal{L}_{decoder}=\underset{\begin{subarray}{c}\epsilon\sim N(0,I)\\ t\sim[0,T],\\ (z_{x},~{}z_{y})\end{subarray}}{\mathbb{E}}\Big{[}||\epsilon-h_{\phi}(x^{(t)},t,z_{x},z_{y})||_{2}^{2}\Big{]}.$ (4)
 For models like Kandinsky v2.2, we drop the $z_{y}$ to condition the model on $z_{x}$. Therefore, $\lambda$-ECLIPSE also only conditions the image generation with $z_{y}$ in the prior stage.
@@
 Figure 8: Examples of image-text interleaved training data.
-Left column shows the input the prior model and right images shows the ground truth corresponding images.
+The left column shows the input of the prior model and the right images shows the ground truth corresponding images.
 Note: these examples are generated from $\lambda$-ECLIPSE for better interpretability.
@@
 We partitioned our dataset into an 80:20 split between training and validation, reserving the remaining 1.6 million images for training and the rest for validation.
+Appendix C Implementation Details
+The $\lambda$-ECLIPSE transformer prior architecture is significantly more compact compared to other Text-to-Image (T2I) methodologies.
+Our model employs a configuration of 16 Attention Heads, each with a dimension size of 32, alongside a total of 10 layers.
+Additionally, the embedding dimension size for our model is set at 1280, supplemented by 4 auxiliary embeddings (including, one for canny edge map).
+As $\lambda$-ECLIPSE is not a diffusion prior model, we do not keep time embedding layers.
+Overall, the $\lambda$-ECLIPSE model comprises approximately 34 million parameters, establishing it as a streamlined yet effective solution for Personalaized-T2I.
+Notably, the standard UnCLIP T2I priors contain 1 billion parameters.
+Appendix D $\lambda$-ECLIPSE with Finetuning
+Figure 9: DreamBooth (Stable Diffusion v1.5) vs. $\lambda$-ECLIPSE  (with fine-tuning) w.r.t. DINO and CLIP-T metrics on Dreambench.
+As demonstrated in the main paper (Table 2), the superiority of fine-tuning-based personalization methodologies, whether applied to single-subject or multi-subject frameworks, over non-fine-tuning alternatives is evident.
+Consequently, we have expanded our analysis through additional fine-tuning of the $\lambda$-ECLIPSE.
+Experimental Setup.
+Given that $\lambda$-ECLIPSE effectively trains the T2I prior, capturing concept-specific features to a significant degree, we opted not to further optimize this component.
+Our focus shifted to exclusively fine-tuning the diffusion UNet model ($h_{\phi}$), employing the AdamW optimizer at a learning rate of 1e-5, without warm-up steps.
+For the DreamBooth application within the Stable Diffusion v1.5 model, we selected a learning rate of 5e-6, maintaining consistency in other hyperparameters.
+To simplify, we excluded the use of a prior preservation regularizer and conducted training on the Dreambench platform using a single RTX A6000 GPU.
+Results.
+Our findings, illustrated in Figure 9, reveal that $\lambda$-ECLIPSE and DreamBooth exhibit improved performance with incremental fine-tuning steps.
+Notably, the DINO score improved from 0.61 to 0.68 with few optimization steps and outperforms the baselines (see Table 2).
+A detailed analysis indicates that while DreamBooth’s DINO score improves, its CLIP-T performance diminishes, hinting at concept overfitting.
+Conversely, $\lambda$-ECLIPSE consistently improves in DINO scoring without adversely impacting the CLIP-T performance, underscoring the efficacy of our image-text interleaved training approach at the prior stage.
+Qualitative comparisons, as shown in Figure 10, further highlight the benefits of fine-tuning $\lambda$-ECLIPSE with minimal steps.
+Advantages of fine-tuning $\lambda$-ECLIPSE.
+The fine-tuning of $\lambda$-ECLIPSE, in comparison to the baselines, reveals two key benefits: 1) Achieving state-of-the-art (SOTA) performance within a few finetuning steps.
+2) Unlike the Stable Diffusion model, which exhibits catastrophic forgetting of nearby concepts post-DreamBooth fine-tuning, $\lambda$-ECLIPSE maintains previous knowledge.
+This suggests that a single model is sufficient to effectively fine-tune across multiple concepts together.
+This analysis underscores the strategic advantages and enhanced efficiency of fine-tuning $\lambda$-ECLIPSE for personalized applications in complex visual data processing.
+Appendix E Extended P-T2I Baselines Comparison
 Table 7: The detailed overview of subject-driven text-to-image generative methodologies. * represents the backbone base models listed are subject to potential updates or modifications.
-Method Multi-Concept Finetuning Free Base Model Image Input
+Method Multi-Subject Finetuning-Free Base-Model # of Input Images
 Re-Imagen [9] Imagen Single
@@
 SSR-Encoder [54] SDv1.5 Single
+Emu2 [46] SDXL Single
 […]
-Appendix C Implementation Details
-The $\lambda$-ECLIPSE transformer prior architecture is significantly more compact compared to other Text-to-Image (T2I) methodologies.
-Our model employs a configuration of 16 Attention Heads, each with a dimension size of 32, alongside a total of 10 layers.
-Additionally, the embedding dimension size for our model is set at 1280, supplemented by 4 auxiliary embeddings (including, one for canny edge map).
-As $\lambda$-ECLIPSE is not a diffusion prior model, we do not keep time embedding layers.
-Overall, the $\lambda$-ECLIPSE model comprises approximately 34 million parameters, establishing it as a streamlined yet effective solution for Personalaized-T2I.
-Notably, the standard UnCLIP T2I priors contain 1 billion parameters.
-Appendix D Extended P-T2I Baselines Comparison
-We further expand our comparative analysis of P-T2I methods encompassing a total of 32 approaches including ours and parallel works.
-Table 5 summarizes them into four crucial aspects: 1) multi-concept support, 2) fine-tuning free, 3) base model types, and 4) the required number of input images.
-To summarize, $\lambda$-ECLIPSE is the only methodology built on top of the UnCLIP models while supporting multi-concept personalization, fine-tuning free, and only requires a single reference concept image.
+We further expand our comparative analysis of P-T2I methods encompassing a total of 33 approaches including ours and parallel works.
+Table 7 summarizes them into four crucial aspects: 1) multi-subject support, 2) fine-tuning free, 3) base model types, and 4) the required number of input images.
+To summarize, $\lambda$-ECLIPSE is the only methodology built on top of the UnCLIP models while supporting multi-subject driven image generation with fine-tuning free, and only requires a single reference image input for the training.
 We detail the comparison below:
-Multi-Concept Generation.
-Multi-concept generation enables users to integrate multiple personal subjects to generate an image that follows the text prompts and aligns with all the concept visuals.
-Of the 32 methods analyzed, 14 offer this capability.
-Ten methods require separate training for each concept and then an additional fusing step for combining the learned concepts.
-(Mix-of-Show).
-While only four methods support fast multi-concept personalization.
-Only a few can learn auxiliary guided information such as canny edge, depth maps, or open-pose and adapt style variation.
+Multi-Subject Generation.
+Multi-subject generation enables users to integrate multiple personal subjects to generate an image that follows the text prompts and aligns with all the concept visuals.
+In total, 15 of the 33 methods offer this capability, while 6 methods support fast multi-subject personalization, others demand separate training for each subject to be learned and then an additional fusing step for combining the learned subjects is required (i.e.
+Zip-LoRA, Mix-of-Show).
+Among these methods, only a few can learn auxiliary guided information such as canny edge, depth maps, or open-pose and adapt style variation (i.e.
+Kosmos-G).
 Fine-tuning Free (Fast Personalization).
-Many models require test-time fine-tuning.
+Many methods require test-time fine-tuning.
 Each varies on which part alteration occurs, as early models tend to modify the whole UNet.
 In contrast, recent models tune a small portion of the cross-attention layers or introduce additional layers performing as adapters.
-In our analysis of P-T2I models, 13 out of 32 models employ a fine-tuning free approach which enables fast personalization.
+In our analysis of P-T2I methodologies, 14 out of 33 methods employ a finetuning-free approach which enables fast personalization.
 Diffusion Independent.
 A majority of the reviewed models utilize diffusion models, with Stable Diffusion being the predominant choice, spanning versions 1.4, 1.5, 2.1, and XL.
-Few adapt Imagen (SuTi, Taming) and interestingly Mix-of-show employs ChillOutMix as their pre-trained model, known for its adeptness at preserving realistic concepts like human faces.
+Few adapt Imagen (SuTi, Taming) and Mix-of-show employs ChillOutMix as their pre-trained model, known for its adeptness at preserving realistic concepts like human faces.
 A unique outlier in this landscape is our $\lambda$-ECLIPSE, the only one that eschews the use of any diffusion prior model.
-Easiness to Use (Input and Output).
-Ease of use, particularly regarding model input and output, significantly impacts user convenience in subject-driven generalization.
-A more user-friendly model typically requires a single input image, as opposed to multiple images of the same subject.
-In our study, 18 models offer P-T2I capabilities with just one input image.
+Easiness of Use.
+A more user-friendly model typically requires a single reference image per subject, as opposed to multiple images of the same subject.
+In our study, 19 methods offer P-T2I capabilities with just one input image.
 In contrast, others often require 4 to 5 images of the subject.
-Additionally, some models necessitate storage space for learned concepts, ranging from a few hundred kilobytes (e.g., Perfusion, HyperDreamBooth) to several megabytes (e.g., Zip-LoRA).
-Our model stands out by eliminating the need for individual concept pre-learning or storing any artifacts for P-T2I utilization, offering a streamlined, efficient user experience.
-Appendix E More Qualitative Results
-In this section, we present an array of extended examples of P-T2I generation, which presents complex compositions that are hard to achieve in $\lambda$-ECLIPSE as well as other models.
-Figure 10 showcases the complexity level that notably escalates as it goes from top to bottom.
-As the complexity of the visual concepts increases, we observe that all methodologies perform poorly in maintaining the subject-fidelity.
-This applies to $\lambda$-ECLIPSE as well.
-Interestingly, $\lambda$-ECLIPSE retains the compositional understanding while baselines perform poorly in all scenarios.
-In addition, we demonstrate some examples of how P-T2I methods exhibit inconsistency in their generated results depending on each trial.
-Figure 11 exhibits consistency in single and multiple concept generation across both models.
-However, Kosmos-G exhibits inconsistency in generating multiple concepts, sometimes placing the Ironman suit in various parts of the dog or omitting it entirely.
-This also means that $\lambda$-ECLIPSE reduces the diversity of the generated images to improve the consistency.
-However, this behavior is observed across the UnCLIP family of models.
-In our final exploration, we examine the preservation of actual human facial characteristics in target scenarios, combined with varying captions.
-While each model attempts to retain the original face, none can precisely replicate the exact personal facial features.
-These examples generally fail to accurately represent the intended compositions, except in one instance in IP-Adapter FaceID.
-Appendix F Limitations
-Our work represents a pioneering effort in harnessing a method that utilizes the latent space of pre-trained CLIP models for P-T2I generation, it is important to acknowledge certain limitations.
-Firstly, CLIP, despite its capabilities, does not perfectly capture hierarchical representations, and this limitation can occasionally result in suboptimal outcomes.
-Consequently, the CLIP contrastive loss can cause the model to deviate from the original subject features, making performing P-T2I on complex concepts, such as human faces, more challenging.
-We anticipate that the enhancement of CLIP representations will correspondingly improve the performance of our framework in the realm of P-T2I mapping.
-The $\lambda$-ECLIPSE model has been trained with 34 million parameters and 1.6 million images.
-However, increasing the data quality and parameters might lead to improvements in the results.
-Appendix G Broader Impact
-The recent retraction of the LAION-5B dataset, prompted by concerns over the presence of data samples potentially breaching ethical guidelines, has significantly affected various research initiatives.
-This situation underscores the crucial importance of careful dataset selection and raises a cautionary note for users about the ethical dimensions involved.
-In line with this, the community needs caution in the application of associated methodologies as well.
+Additionally, some methods necessitate storage space for learned concepts, ranging from a few hundred kilobytes (e.g., Perfusion, HyperDreamBooth) to several megabytes (e.g., Zip-LoRA).
+Our method stands out by eliminating the need for individual concept pre-learning or storing any artifacts for P-T2I utilization, offering a streamlined, efficient user experience.
+Appendix F Multibench Dataset
+Table 8: Example of prompt templates used for Multibench dataset. Subjects presented in Table 9 are placed in {}.
+Two subjects Three subjects
+{} in the {} {} with a {} and {}
+{} wearing a {} {} is playing with {} in {}
+{} chasing a {} {} with {} in front of {}
+{} looking at a {} {} with a {} and a view of the {}
+{} is sitting on a {} {} with a {} and {} in the background
+{} standing on a {}
+{} and {} playing in the garden
+{} and {} on top of the mountain
+{} and {} in the jungle
+{} and {} in the snow
+{} and {} on the beach
+{} and {} on a cobblestone street
+{} and {} standing next to each other
+Table 9: Number of occurrences of unique subject categories.
+The left side of the table are subjects used for two subjects prompts, and the right side of the table are subjects used for three subjects prompts.
+Two subjects Three subjects
+dog 76 boat 5 dog 81 rainbow 35
+cat 76 park 4 stuffed animal 105 ruins 35
+bird 76 ruins 9 toy 105 tower 35
+horse 73 castle 5 cat 81 horse 81
+guinea pig 73 desert 4 desert 60 bird 81
+glasses 5 rainbow 5 hill 60 guinea pig 81
+hat 5 candle 5 castle 45 guitar 25
+tower 10 backpack 3 backpack 65 french horn 25
+can 130 vase 25
+candle 65 robot 25
+church 35
+We provide additional qualitative results in Figure 5.
+For the multi-subject image benchmark, our dataset comprises 2,308 unique prompts, segmented into 904 two-subject and 1,476 three-subject prompts.
+This dataset integrates subjects from the original DreamBench dataset, featuring 30 distinct concepts.
+We expanded the dataset by incorporating additional concepts vital for two and three subject-specific prompts, such as various parks, hats, glasses, and more.
+Prompt templates and the count of unique subject categories featured in prompts are detailed in Tables 8 and  9, respectively.
+Overall, the dataset includes 217 two-subject compositions and 405 three-subject compositions, enriching the benchmark’s diversity and comprehensiveness.
+Appendix G Qualitative Results & Failure Cases
+In this section, we showcase a collection of detailed qualitative examples from the P-T2I generation process, highlighting the challenges of crafting complex compositions within $\lambda$-ECLIPSE and comparative models.
+As depicted in Figure 11, the complexity of the showcased examples progressively increases, illustrating a noticeable escalation in the intricacy of visual concepts from the top to the bottom of the figure.
+With the rising complexity, we note a universal decline in the ability of all methodologies, including $\lambda$-ECLIPSE, to preserve subject fidelity accurately.
+Interestingly, despite these challenges, $\lambda$-ECLIPSE demonstrates a better grasp of compositional integrity, unlike the baseline models which falter across all complexity levels.
+Moreover, we present instances demonstrating the variability in outcomes produced by P-T2I methods across different trials.
+As illustrated in Figure 12, while there is a semblance of consistency in generating single and multiple concepts between models, Kosmos-G specifically shows variability in rendering multiple concepts—occasionally misplacing elements of the Ironman suit on a dog or failing to include it altogether.
+This phenomenon suggests that $\lambda$-ECLIPSE minimizes image diversity to enhance result consistency, a trait observed across the UnCLIP model family.
+Figure 10 offers qualitative insights into the performance of $\lambda$-ECLIPSE without and with minimal fine-tuning.
+It is evident that in certain edge cases, where $\lambda$-ECLIPSE initially struggles to fully grasp novel visual concepts without finetuning, a modest application of few optimization iterations significantly enhances concept capture.
+Further optimization not only preserves text composition but also enriches minor, subject-specific details, underscoring the adaptability and finesse of $\lambda$-ECLIPSE in nuanced image generation.
+Moreover, in our evaluations using the Multibench dataset, we noticed that both the baseline models (Kosmos-G and Emu2) and $\lambda$-ECLIPSE encounter difficulties in precisely maintaining all subject-specific details, as depicted in Figure 14.
+This underscores that zero-shot multi-subject P-T2I generation remains a significant challenge in the field.
+Further, we explored how well each model preserves genuine human facial characteristics in various scenarios, particularly when combined with differing captions.
+The qualitative examples in Figure 13 shed light on this aspect.
+Although each model strives to maintain the original facial features, none succeeds in replicating the specific personal facial details accurately.
+These instances typically fall short of precisely conveying the intended compositions, with the exception of one scenario in IP-Adapter FaceID, indicating a notable area for future improvements in model performance.
+Appendix H Limitations
+Our work marks a pioneering venture into leveraging the latent space of pre-trained CLIP models for P-T2I generation.
+Nonetheless, it’s crucial to recognize certain constraints.
+Primarily, despite its strengths, CLIP’s inability to perfectly capture hierarchical representations occasionally leads to less-than-ideal results.
+This issue, stemming from the CLIP contrastive loss, can cause deviations from the original subject features, particularly when generating P-T2I for complex subjects like human faces.
+We believe that enhancing CLIP’s representations could significantly boost our framework’s efficacy in P-T2I mapping.
+The $\lambda$-ECLIPSE model, trained on 34 million parameters and 1.6 million images, presents a substantial foundation.
+Yet, there’s potential for further refinement, as increasing the quality of data and the number of parameters could yield even better outcomes.
+Appendix I Broader Impact
 Subject-driven image generation or Personalized Text-to-Image (P-T2I) methods have the potential to be a transformative tool in numerous domains.
@@
 On the other hand, the ease of altering and creating images raises concerns about the responsible use of this technology which requires significant ethical and legal considerations.
-We recommend developers provide a more secure way (such as image attribution) for end-users to ensure accountability for misuse of such models [18].
 Users must be acutely aware of being able to infringe intellectual property rights and create misleading or harmful content.
+We recommend developers provide a more secure way such as image attribution [18] for end-users to ensure accountability for misuse of such models.
 As such, those employing subject-driven image-generation techniques should exercise careful judgment, ensuring that their work adheres to ethical standards and legal boundaries.
 It is imperative that the broader implications of this technology are considered, and that a commitment to responsible and conscientious use guides its application.
-Figure 10: Qualitative examples of increasing complexity of novel visual concepts as we move from top to bottom.
+Figure 10: Qualitative examples of $\lambda$-ECLIPSE without finetuning and different stages of finetuning.
+Figure 11: Qualitative examples of the increasing complexity of novel visual concepts as we move from top to bottom.
 Figure 12: Qualitative examples of showcasing the consistency comparisons between Kosmos-G and $\lambda$-ECLIPSE.
 Figure 13: Qualitative examples of showcasing the failure cases on human faces on Kosmos-G, IP-Adapter (SDXL), IP-Adapter (FaceID), and $\lambda$-ECLIPSE.
+Figure 14: Qualitative examples of showcasing the failure cases on Multibench of Kosmos-G, Emu2, and $\lambda$-ECLIPSE.